### v1

# : Context Management, Validation, and Transaction Guarantees for Multi-Agent LLM Planning

Abstract.
Recent LLM-based agent frameworks have demonstrated impressive capabilities in task delegation and workflow orchestration, but face significant challenges in maintaining context awareness and ensuring planning consistency. This paper presents , a structured multi-agent framework that addresses four fundamental limitations in current LLM approaches: inadequate self-validation, context narrowing, lacking transaction properties, and insufficient inter-agent coordination. By implementing specialized context management agents and validation protocols, preserves critical constraints and state information throughout complex planning processes, enabling robust and consistent decision-making even during disruptions. We evaluate our approach using selected problems from the REALM benchmark, focusing on sequential and reactive planning scenarios that challenge both context retention and adaptive reasoning. Our experiments with state-of-the-art LLMsβClaude 3.7, DeepSeek R1, GPT-4o, and GPT-o1βdemonstrate that while these models exhibit impressive reasoning capabilities, they struggle with maintaining global constraint awareness during complex planning tasks, particularly when adapting to unexpected changes. In contrast, the distributed cognitive architecture of shows significant improvements in planning consistency, constraint enforcement, and adaptation to disruptions in various scenarios.
1. Introduction
Multi-Agent Systems (MAS) have long been a cornerstone of distributed computing and database systems (Lesser etΒ al., 2004; Jennings, 1993; Dunin-Keplicz and Verbrugge, 2010). Over the past few decades, their development has followed two primary trajectories. In the database community, MAS integrated robust transaction processing principles such as ACID properties (Gray, 1981), ensuring consistency and reliability for complex multi-step operations. In contrast, distributed systems research has prioritized coordination protocols and flexible collaboration mechanisms (Wooldridge, 2009; Durfee, 1999), enabling scalable multi-agent interactions without the constraints of strict locking or heavyweight transactional guarantees. These parallel approaches have resulted in frameworks optimized for different priorities: transactional integrity versus adaptive coordination, highlighting the trade-off between strong consistency and flexible execution in practical systems.
Recent advances in Large Language Models (LLMs) have revitalized MAS as a paradigm for sophisticated reasoning and multi-agent collaboration (Du etΒ al., 2023; Chang, 2023, 2024b). Frameworks such as AutoGen (Wu etΒ al., 2024), LangGraph (LangChain AI, 2024), and CAMEL (Li etΒ al., 2023a) demonstrate how LLM-based agents can decompose tasks, interact across modalities, and coordinate to solve complex problems. However, this resurgence often neglects the foundational transaction guarantees essential to reliable multi-agent workflows, particularly in domains that require robust state management. Unlike traditional MAS, LLM-based systems often lack mechanisms for maintaining strong consistency, failure recovery, and rollback handling, leading to inconsistent states, partial failures, and unreliable execution in real-world applications.
For example, in a travel-booking scenario, an LLM-based MAS may independently issue flight and hotel reservations without ensuring their coordinated success. If the flight is later canceled, the system may fail to recognize the inconsistency, leaving the hotel booking active. While a monolithic LLM could attempt to track such dependencies internally, context loss in long conversations (Modarressi etΒ al., 2025; Xiao etΒ al., 2024; Hsieh etΒ al., 2024; Liu etΒ al., 2024a) can cause it to forget earlier steps, leading to contradictory or incoherent decisions. Moreover, splitting tasks across multiple agents exacerbates the problem, as no built-in supervisory mechanism reconciles their state changes. Compounding this issue, LLMs inherently struggle with self-validation, as highlighted by GΓΆdelβs incompleteness theorems (GΓΆdel, 1967; Chang, 2024a), making them unreliable for detecting and correcting their own errors. These challenges demand a transactional framework that preserves the intelligence and adaptability of LLM-based MAS while ensuring consistency and reliability in long-running, interdependent workflows.
To address these limitations, we propose , a Transactional Multi-Agent System that extends the Saga pattern (Garcia-Molina and Salem, 1987; Weikum and Vossen, 2001)βoriginally designed for long-lived transactions by decomposing complex operations into smaller, compensable units. By integrating transactional logic and rollback mechanisms into LLM-based MAS, ensures that each operation is fully committed or coherently reversed, preventing partial failures and inconsistencies. For example, in a travel-booking scenario, a failed flight reservation would automatically trigger compensatory rollbacks of associated hotel and train bookings, maintaining consistent global states without manual intervention.
To effectively address LLM-specific challenges and context loss problems, introduces three critical innovations that directly support our four primary contributions:
β’
Spatial-Temporal State Identification and Checkpointing provides the foundation for transactional correctness by tracking the evolution of agent interactions and enabling precise state recovery.
β’
Inter-Agent Dependency Management ensures consistency by monitoring interdependencies and alerting agents to unresolved constraints, enhancing both the reliability of the system and intelligent coordination.
β’
Independent Validation of Critical Junctures enhances correctness by employing specialized small-context agents with rigorous validation criteria, effectively addressing the self-validation limitations of individual LLMs.
This hybrid approachβintegrating transactional processing with adaptive multi-agent intelligenceβmakes particularly effective for complex real-world applications in financial trading, healthcare, travel coordination, supply chain management, and emergency response, to name just a few.
Contributions
The primary contributions of this work are as follows:
Transactional Correctness: introduces transactional safeguards to LLM-based MAS, ensuring strong consistency across multi-agent workflows.
Robust Validation: incorporates context management and verification mechanisms to improve execution correctness and prevent inconsistencies.
Intelligence: By integrating LLM-based reasoning with transactional coordination, enables adaptive, explainable, and context-aware planning.
Scalable Performance: minimizes overhead via modularized validation, balancing safety and efficiency for real-world deployments.
2. Related Work
We discuss related work in four parts. First, we examine the evolution of transaction management in distributed systems, particularly focusing on the Saga pattern. Second, we analyze the limitations of LLMs that necessitate our three key requirements: transactional integrity, independent validation, and context preservation. Third, we review current multi-agent LLM frameworks and their transaction-related limitations. Finally, we position in relation to recent approaches that attempt to address similar challenges.
2.1. Transaction Management Systems
Transaction management has evolved significantly since the introduction of ACID properties by Gray (Gray, 1981). In distributed systems, the challenges of maintaining these properties across service boundaries led to alternative models such as BASE (Basically Available, Soft state, Eventually consistent) (Pritchett, 2008) and the development of specialized patterns for long-lived transactions.
The Saga pattern, introduced by Garcia-Molina and Salem (Garcia-Molina and Salem, 1987), addresses the limitations of traditional ACID transactions in distributed, long-running processes. Unlike atomic transactions that rely on locks, Sagas break complex operations into sequences of smaller, local transactions, each with a corresponding compensating transaction that semantically undoes its effects. This approach has become fundamental in micro-service architectures and event-driven systems (Richardson, 2018), allowing them to maintain logical consistency without the performance penalties of distributed locking.
Workflow management systems such as YAWL (vanΒ der Aalst and ter Hofstede, 2005) and more recent implementations on cloud platforms (aws, 2023; azu, 2023) have incorporated Saga-like patterns to manage complex business processes. However, these systems typically rely on predefined, rigid workflows with limited adaptability to unforeseen circumstancesβa limitation that becomes particularly acute when integrating with AI systems that must reason about unexpected events or novel scenarios.
The fundamental principles of transaction management established in these systemsβatomicity, compensation-based recovery, and state consistencyβprovide a theoretical foundation for , which extends these concepts to the unique challenges of LLM-based agent workflows.
2.2. LLM Limitations Necessitating βs Key Requirements
Our analysis of current LLMs reveals three fundamental limitations that directly motivate the core requirements of : the need for transactional integrity, independent validation, and strategic context preservation.
Self-Validation Gap Necessitating Independent Validation
LLMs lack mechanisms for rigorous self-validationβa limitation directly connected to GΓΆdelβs insights about formal systems (GΓΆdel, 1967). Recent studies confirm that while self-refinement methods (Madaan and Yazdanbakhsh, 2022; Li etΒ al., 2023b; Jiang etΒ al., 2024) iteratively improve reasoning using the LLMβs own feedback, they struggle to correct errors beyond the modelβs inherent capability ceiling (Huang etΒ al., 2024). In transactional contexts, this limitation manifests itself as factual inconsistencies that can lead to invalid operations, the inability to verify the feasibility of the plan, and the difficulty in maintaining consistency between operations (Yamin etΒ al., 2024). These challenges directly motivate βs independent validation framework.
Statelessness Necessitating Transactional Integrity
LLMs process each prompt independently, lacking inherent mechanisms to maintain the state between interactions. This fundamental statelessness makes transaction management essential, as there are no built-in guarantees that a sequence of operations will be treated as a cohesive unit or that failure in one operation will trigger appropriate compensations in others. Without transaction management, LLM-based systems risk inconsistent states, lost operations, and the inability to recover coherently from failures.
Context Limitations Necessitating Strategic Preservation
The self-attention mechanism of LLMs gives higher weight to immediate context, causing a narrowing of focus over longer sequences. Empirical measurements reveal that context retention significantly degrades beyond certain token thresholds (Xiao etΒ al., 2024; Modarressi etΒ al., 2025), and the accuracy of information recall dramatically decreases for content in middle segments of long contexts (Liu etΒ al., 2024a; Hsieh etΒ al., 2024). These limitations directly motivate β strategic context preservation mechanisms, as the system must explicitly track and maintain critical information that LLMs might otherwise βforgetβ during long-running transactions.
Together, these limitations create a compelling case for addressing all three key requirements of β. Without comprehensive transaction management, independent validation, and strategic context preservation, LLM-based agent systems cannot provide the reliability guarantees needed for critical real-world applications.
2.3. Multi-Agent LLM Frameworks and Their Transaction Limitations
Recent frameworks such as AutoGen (Wu etΒ al., 2024), LangGraph (LangChain AI, 2024), CAMEL (Li etΒ al., 2023a), and MACI (Chang, 2024a) extend multi-agent principles to LLM-based coordination but fall short in addressing the three key requirements that targets.
Transactional Integrity Gaps
LangGraph provides basic state management and workflow orchestration but does not implement comprehensive transaction support. While its graph-based workflow model allows for structured agent interactions, it lacks built-in mechanisms for atomicity guarantees, compensation handling, or failure recovery across complex multi-step operations. AutoGen similarly focuses on flexible agent communication but provides no systematic approach to rolling back previous actions when operations fail mid-workflow, potentially leaving the system in inconsistent states.
Validation Limitations
Current frameworks typically rely on the LLMβs inherent capabilities for validation, without the independent verification mechanisms that implements. This reliance on self-validation makes these systems vulnerable to the very limitations discussed in Section 2.2, as they lack robust safeguards against LLM errors, hallucinations, or inconsistencies between agent boundaries.
Context Preservation Challenges
While frameworks like CAMEL maintain conversation history, they provide minimal support for strategic context preservation across long-running operations. Their approaches to context management focus on communication rather than the systematic tracking of operation history, dependencies, and state changes needed for reliable transaction processing.
These frameworks serve primarily as integration platforms rather than addressing the deeper challenges related to transaction management, independent validation, and context preservation. Their focus on orchestration without these critical capabilities creates the gap that addresses.
2.4. Transaction-Oriented Approaches in LLMs
Recent approaches have begun to address aspects of transaction management in LLM-based systems, though none provides the comprehensive solution that offers.
Transactional Enhancements to LLM Frameworks
Recent work on enhancing MAS with transactional capabilities (Gao etΒ al., 2024) represents an important step toward more reliable LLM workflows. However, these improvements focus primarily on state management without comprehensive compensation mechanisms, independent validation, or preservation of the strategic context provided by . Similarly, experimental transaction support in other frameworks often treats transactions as an add-on rather than a foundational component integrated throughout the system.
Reliability-Focused Planning Systems
Systems like PLASMA (Brahman etΒ al., 2024) address reliability in planning but lack transaction-specific mechanisms like compensation handling. Similarly, search-based approaches such as LLM-MCTS (Zhao etΒ al., 2023) and Tree-of-Thought methods (Yao etΒ al., 2024) improve plan generation but provide limited support for execution guarantees or recovery when operations fail. These systems typically focus on pre-execution planning rather than the runtime transaction management needed for maintaining consistency through a workflowβs lifecycle.
Agent Coordination with Limited Transaction Support
Systems like ADAS and AFlow (Hu etΒ al., 2024; Zhang etΒ al., 2024) implement agent coordination protocols but lack robust transaction management. Their approaches to handling failures are typically ad hoc rather than following established transaction patterns like the Saga approach. This gap becomes particularly apparent when dealing with partial failures in complex workflows, where these systems struggle to maintain consistency across agent boundaries.
advances beyond these approaches by providing a comprehensive framework that integrates transaction management principles, independent validation mechanisms, and strategic context preservation. By addressing all three key requirements simultaneously, enables LLM-based multi-agent systems to achieve reliability guarantees.
3. System Requirements for
Building on the foundational concepts introduced in SectionΒ 1, we now formally define the requirements for . Our framework extends beyond the mere orchestration capabilities of current LLM-based systems to address three critical needs: transactional integrity, independent validation, and context filtering and preservation.
The strategic context filtering and preservation requirement addresses a fundamental challenge with LLM agents: balancing comprehensive context retention against the practical limitations of context windows and attention mechanisms. By intelligently filtering which information is preserved across operations, ensures that critical data for transaction integrity is maintained while avoiding context overload.
3.1. Transactional Requirements
should provide transactional guarantees (Gray, 1981) that extend beyond individual agent operations to encompass entire workflows when strong consistency is required. For workflows requiring such guarantees, the system ensures atomicity, treating a sequence of operations performed by multiple agents as an indivisible unit. Given a system state , applying to results in either state (all operations applied) or state (no operations applied), with no intermediate states persisting. However, certain multi-agent workflows may operate effectively with eventual consistency or other relaxed guarantees when appropriate for their domain requirements.
Consistency requirements specify that the system should transition from one valid state to another. Given a set of invariants , if satisfies , then should also satisfy . This is particularly challenging when invariants span multiple agentsβ knowledge domains.
For isolation, concurrent workflows should not interfere with each other. Any two concurrent transaction sequences and should result in a state equivalent to either applying followed by , or followed by . Durability aims to ensure that once a workflow is committed, its effects persist despite system failures, with all state changes retrievable after system restart.
Furthermore, while several transaction management patterns could be applied to LLM agent workflows, adopts the Saga pattern as defined by Garcia-Molina and Salem (Garcia-Molina and Salem, 1987) due to its particular suitability for long-running, distributed operations involving autonomous agents that may require compensating actions. maps each operation to a transaction with a corresponding compensating transaction , forming:
where each is a local transaction and each is its corresponding compensating transaction. If any fails, all previously executed transactions are compensated by executing in reverse order.
Alternative approaches such as two-phase distributed commit protocols (Bernstein etΒ al., 1987; Gray and Reuter, 1993) could provide stronger consistency guarantees, but would impose higher coordination overhead and potentially block agent operations during uncertainty periods. Similarly, eventual consistency models (Pritchett, 2008) might offer greater scalability but would sacrifice the clear semantic guarantees needed for critical workflows. The Saga pattern strikes an effective balance, providing meaningful transactional guarantees while accommodating the distributed and long-running nature of LLM agent operations.
| Mechanism | Information Recorded |
| Application State () | |
| Domain Entities | β’ Application-domain data objects β’ Domain entity states β’ State checkpoints and snapshots |
| Operation State () | |
| Execution Records | β’ Operation inputs and outputs β’ Timestamps and execution metrics β’ Operation completion status |
| Decision Reasoning | β’ LLM reasoning chains β’ Decision justifications β’ Alternative options considered |
| Compensation Actions | β’ Semantic inverse operations β’ Compensation prerequisites β’ Recovery state requirements |
| Dependency State () | |
| Causal Dependencies | β’ Inter-operation relationships β’ Data flow dependencies β’ Resource allocation linkages |
| Constraint Satisfaction | β’ Boolean condition evaluations β’ Satisfaction evidence β’ Decision timestamps |
3.2. Transaction State Management
addresses the challenge of enforcing ACID guarantees in a distributed LLM agent environment with stateless agents. implements spatial-temporal state identification through a comprehensive transaction state management approach that preserves selective context information throughout multi-agent workflows. Using the MACI algorithm (Chang, 2024a) that we have previously developed, the states and dependencies of a given application specification can be automatically extracted to enable effective transaction state management.
3.2.1. State Dimensions and Context Preservation
maintains three critical state dimensions to support both transaction management and context preservation:
β’
Application State (): The application-domain dependent data being manipulated (e.g., travel booking details).
β’
Operation State (): Records of executed operations, including inputs, outputs, and effects on domain entities.
β’
Dependency State (): An explicit mapping of inter-operation dependencies (constraints) and their satisfaction criteria.
By structuring operations, dependencies, and application data, ensures that essential information remains accessible even as LLMsβ attention mechanisms might otherwise cause older context to be lost. TableΒ 1 details the specific information recorded and mechanisms used to implement these state dimensions. Note that the decision reasoning recorded under is a critical component where LLMs make significant contributions that traditional workflow engines cannot support.
For example, in the Thanksgiving dinner planning scenario from SectionΒ 5.2, would maintain application state for each family memberβs location and activities, operation state for travel arrangements and food preparation steps, and dependency state capturing constraints like βturkey must be in the oven by 2:00 PMβ and βJames must pick up Emily after exiting the terminal.β
3.2.2. Dependency Tracking for Compensation
maintains formal representations of operation dependencies in a directed graph structure:
where represents the condition that must be satisfied for the dependency to be resolved. These conditions can be complex boolean expressions combining multiple upstream requirements, such as:
where represents an arbitrary Boolean function that determines how the upstream conditions () are combined.
This dependency graph is crucial for compensation planning during failures. When an operation fails, traverses the dependency graph to identify the operations that may be affected, creating a precise compensation sequence that ensures that the consistency of the system is maintained. The graph also enables to identify and preserve only the critical dependencies needed for potential compensation, reducing overhead.
In the wedding coordination example from SectionΒ 5.3, when the traffic alert is received, the dependency graph allows to precisely determine which segments of Patβs journey are affected and what downstream activities (gift shop, tailor visit) need to be replanned or compensated.
3.2.3. Intra-Agent Critical Context Tracking
preserves critical context that LLMs might otherwise forget during extended operations. This includes agent specifications, intermediate reasoning states (chain-of-thought), decision justifications, and the context required for potential compensation actions. Empirical study (Xiao etΒ al., 2024; Modarressi etΒ al., 2025; Liu etΒ al., 2024a; Hsieh etΒ al., 2024) shows that LLMs can forget simple facts explicitly provided in context, such as travel times between locations.
The system implements a validation cycle where agent operations produce selected logged context that undergoes validation against system specifications. Valid operations proceed to the next workflow node, while invalid ones trigger compensating actions before retrying. This ensures that:
β’
Compensation mechanisms have necessary historical context.
β’
Operations remain consistent with defined specifications.
β’
Dependency relationships maintain integrity throughout the workflow.
This context tracking directly addresses the context narrowing issue observed in our experiments in SectionΒ 5.2, where models like Claude 3.7 lost track of critical constraints such as fire safety and travel times during reactive planning.
3.2.4. Inter-Agent Communication Protocol
implements a structured communication protocol between agents that explicitly manages what information is passed between them. This protocol ensures that:
β’
Each transaction agent receives precisely the information needed for its operation.
β’
Dependencies and constraints are properly communicated.
β’
boundaries are respected in multi-agent operations.
The protocol explicitly separates the agentβs working context (input prompt, reasoning space) from the transaction state. This separation allows agents to operate with minimal context while the system maintains complete transaction information, addressing scalability concerns in complex multi-agent workflows.
3.3. Independent Validation Framework
With transaction state management ensuring persistence, we now examine how implements its independent validation of critical junctures to address the LLM self-validation limitations described in SectionΒ 1.
3.3.1. Multi-Level Validation Architecture
implements a two-tier validation approach that separates intra-agent and inter-agent concerns.
Intra-Agent Validation
This validation operates within an individual LLM agent, addressing both theoretical limitations (GΓΆdelβs incompleteness) and practical challenges (attention narrowing, context forgetting).
Syntactic Validation: Ensures outputs conform to expected formats, structures, and schemas (e.g., JSON validity).
Semantic Validation: Verifies logical consistency, coherence, and significance of the agentβs operations.
Factual Validation: Confirms that critical facts provided in context are maintained throughout the reasoning process (e.g., travel times, constraints).
Constraint Adherence: Verifies that agent decisions respect all explicitly defined constraints from specifications.
Reasoning Validation: Evaluates whether the agentβs chain-of-thought process follows logically sound inference patterns.
Context Retention: Confirms the agent has maintained necessary contextual information required for its current operation.
In the Thanksgiving dinner scenario, this validation would catch the fire safety violation where Sarah was scheduled to leave while the turkey was unattended in the oven.
Inter-Agent Validation
This validation ensures proper information flow and dependency management across multiple agents in a workflow.
Contract Validation: Verifies that inputs and outputs between agents conform to predefined interfaces.
Dependency Satisfaction: Confirms that all prerequisites are met before an agent operation is executed.
Consistency Checks: Ensures shared data elements maintain consistent meanings across agent boundaries.
Temporal Validation: Verifies that operations happen in the correct sequence, respecting causal dependencies.
Mutual Agreement: For collaborative decisions, confirms that multiple agents have reached consensus on shared facts.
Transaction Boundary Integrity: Ensures that transaction ACID properties are maintained across agent operations.
For example, in the wedding coordination scenario, inter-agent validation would ensure that Patβs position is consistently tracked across agents, preventing the error where DeepSeek R1 βrewrote historyβ by starting Patβs drive to the airport after the traffic alert despite Pat already being at the airport.
3.3.2. Validation Protocols and Response Mechanisms
To operationalize the validation types in TableΒ 2, implements structured protocols that govern system responses to validation outcomes:
Rejection Protocol: Triggers compensating transactions upon validation failure (e.g., canceling hotel bookings if flight validation fails).
Augmentation Protocol: Refines agent outputs while preserving agent autonomy (e.g., formatting restaurant recommendations without changing selections).
Feedback Protocol: Captures validation results to improve future agent operations (e.g., noting that travel time estimates were consistently underestimated).
These protocols directly connect validation to transaction management, ensuring that validation failures trigger appropriate compensating actions while maintaining workflow integrity across multiple agents.
3.4. Failure Handling and Recovery
βs failure handling leverages the transaction state and dependency tracking described in the previous sections to implement comprehensive recovery:
Operation-Level Recovery: When individual operations fail, compensating actions execute to semantically undo their effects.
Workflow-Level Recovery: When entire workflows fail, the system orchestrates multiple compensations in the correct order based on the captured dependency graph.
For each compensating transaction that reverses a failed transaction , the system leverages the preserved transaction state to provide the original inputs and context that received, the outputs and side effects produced, and the specific aspects of that must be undone. This approach ensures reliable compensation even when the original agent is no longer available or when the operation occurred many steps earlier in a complex workflow.
4. Design with Travel Planning
FigureΒ 1 depicts the architecture, which sits between the application layer and LLM multi-agent systems like LangGraph. comprises three frameworks: context management, validation, and transaction. To illustrate the design, we use a travel planning example.
Travel Planning Specification
This application helps users plan international trips with multiple destinations and manage complex itineraries within a specified budget. The system leverages LLMs to generate and refine feasible travel plans, followed by a transactional booking process implemented with to ensure consistency.
4.1. User Requirements
β’
Plan a trip from San Francisco to Berlin and Cologne and then back to San Francisco.
β’
Travel period: June 2025 (flexible within the month)
β’
Budget constraint: $5,000 total.
β’
Required bookings: flights, hotels, and trains between cities.
β’
Preferences:
Moderately priced accommodations (3-4 star hotels).
Direct flights when possible.
Flexible scheduling with 4 days in Berlin and 2 in Cologne.
4.2. System Workflow
The application workflow is divided into two main phases: itinerary planning and booking execution. FigureΒ 2 depicts the work flow with phase one in gray and phase two in green color.
4.2.1. Phase 1: Itinerary Planning
1.
Initial Plan Generation:
1.1.
LLM generates multiple feasible itineraries based on user requirements
1.2.
Each itinerary includes flight options, hotel suggestions, and transportation (train) between cities
1.3.
System estimates costs for each component and validates against budget constraints
1.1.
2.
Iterative Refinement:
2.1.
User reviews suggested itineraries and provides feedback
2.2.
LLM adjusts plans based on feedback while maintaining budget constraints
2.3.
System tracks and preserves critical context (travel dates, preferences, constraints)
2.4.
Process repeats until user is satisfied
2.1.
3.
Plan Finalization:
3.1.
User selects a final itinerary
3.2.
System generates a detailed breakdown of all necessary bookings
3.3.
Critical dependencies between bookings are identified (e.g., hotel check-in must be after flight arrival)
3.1.
4.2.2. Phase 2: Booking Process ( Implementation)
1.
Booking Workflow Creation:
1.1.
System constructs a SAGA transaction sequence based on the finalized itinerary
1.2.
Defines compensation actions for each booking step
1.3.
Establishes validation rules for each transaction
1.1.
2.
Execution:
2.1.
International Flight Booking: SFO Berlin
2.2.
Berlin Hotel Booking
2.3.
Train Booking: Berlin Cologne
2.4.
Cologne Hotel Booking
2.5.
International Flight Booking: Cologne SFO
2.1.
3.
Exception Handling:
3.1.
Potential failures:
Flight unavailability or price changes
Hotel rooms no longer available
Train scheduling conflicts
3.2.
Compensation strategies:
Cancel previous bookings (with cancellation policies)
Replan affected portions of the itinerary
Suggest alternatives within similar price range
3.1.
4.3. Agents, Transactions, and Compensations
FigureΒ 3 presents the organization of the code implementation. The code is organized into three categories: application interface, core, and LangGraph integration. For each agent, we specify its critical context that requires external logging and its validation protocols.
4.3.1. Travel-Planning Agent Specifications
1.
FlightBookingAgent: Handles flight reservations
Critical context: Dates, budget, airline preferences
Validation: Price verification, schedule feasibility, connection times
2.
HotelBookingAgent: Manages accommodation reservations
Critical context: Dates, location constraints, amenity preferences, stars/ratings
Validation: Price verification, availability
3.
TrainBookingAgent: Books rail transportation between cities
Critical context: Travel times, connection with flight arrivals/departures
Validation: Schedule coordination with other bookings
4.
BudgetTrackingAgent: Monitors expenditure
Critical context: Running total of expenses, component costs
Validation: Ensures total remains within budget constraints
5.
ItineraryPlanningAgent: Suggests and refines travel plans
Critical context: User preferences, previous feedback
Validation: Coherence of overall plan, timing feasibility
4.3.2. Transactions and Compensations
1.
Flight Booking Transaction:
Operation: Reserve flight with selected airline
Compensation: Cancel reservation with applicable fees
Validation: Confirm price, schedule, and availability
2.
Hotel Booking Transaction:
Operation: Reserve room for specified dates
Compensation: Cancel reservation according to hotel policy
Validation: Confirm price, dates, and room type availability
3.
Train Booking Transaction:
Operation: Reserve train tickets
Compensation: Cancel tickets according to railway policy
Validation: Confirm the alignment of the schedule with other travel components
4.3.3. Critical Context Tracking
The critical context must be saved so that, upon failure, the compensation or rollback operation can be properly executed. The context is required for undo, but also for rebooking. For instance, if a flight is delayed, the hotel booking can either be changed, or if there is no availability for a new duration, a different hotel can be booked with the saved context.
tracks a list of critical context throughout the travel planning and booking process, which includes travel dates for each city, booking confirmation numbers and references, cancellation policies and deadlines, price information and total budget running, dependencies information between bookings, user preferences and constraints, and travel times between key locations.
The key architectural principles remain the same regardless of the specific details of the itinerary, making this implementation adaptable to various travel planning scenarios. The validation framework provides a foundation for maintaining consistency across different travel configurations.
4.4. Validation Protocols
We have defined for the travel planning example a workflow and five agents. TableΒ 2 further depicts these agentsβ intra-agent validation protocols (ensuring individual agent correctness) and inter-agent validation protocols (maintaining workflow integrity across agents). Intra-agent validation addresses LLM-specific challenges like context forgetting and reasoning errors, while inter-agent validation ensures proper coordination in the distributed workflow. Each validation type is critical for maintaining the reliability of the travel planning system.
| Validation Type | Travel Planning Example |
| Intra-Agent Validation | |
| Syntactic Validation | FlightBookingAgent produces properly structured JSON with all required fields (departure time, arrival time, flight number). |
| Semantic Validation | HotelBookingAgent selects accommodations that cover the entire trip duration without gaps. |
| Factual Validation | TrainBookingAgent maintains the 45-minute travel time from hotel to the train station. |
| Constraint Adherence | BudgetTrackingAgent keeps total trip cost under the userβs maximum budget of $5,000. |
| Reasoning Validation | ItineraryPlanningAgent correctly considers weather forecasts when recommending indoor vs. outdoor activities. |
| Inter-Agent Validation | |
| Dependency Satisfaction | Flights are booked before hotel reservations are finalized, as hotel dates depend on flight availability. |
| Consistency Checks | Location data is represented in the same format (coordinates vs. addresses) across all agents. |
| Temporal Validation | BudgetTrackingAgent completes transactions only after all bookings have been verified and approved. |
| Mutual Agreement | Transportation Agents and ItineraryPlanningAgent agree on feasible travel times between attraction venues. |
| Transaction Boundary Integrity | If flight booking fails, all dependent hotel and train reservations are automatically canceled through compensating transactions. |
For each validation failure, the system implements specific response protocols:
β’
Rejection Protocol: When validation fails critically (e.g., a flight is no longer available), the system triggers compensation for any already-completed transactions.
β’
Augmentation Protocol: For minor validation issues (e.g., missing non-critical information), the system attempts to correct the output without requiring a full retry.
β’
Feedback Protocol: All validation results are logged and used to improve future agent operations, creating a learning system that becomes more reliable over time.
5. Experiments
We design experiments to remedy four shortcomings mentioned in SectionΒ 1 of the current multi-LLM agent systems, namely inadequate self-validation, context narrowing, lacking transaction properties, and insufficient inter-agent coordination.
5.1. Experimental Design
We selected test cases from the REALM benchmark (Geng and Chang, 2025), which evaluates multi-agent systems on eleven distinct problems. For our experiments, we focused on two medium-tier sequential planning challenges (problems #5 and #6) and two reactive planning challenges (problems #8 and #9).
We evaluated four LLMsβClaudeΒ 3.7, DeepSeekΒ R1, GPT-4o, and GPT-o1βalongside our proposed framework. All experiments were conducted between MarchΒ 12 andΒ 17, 2025.
Objective: Coordinate family arrivals and dinner preparation for 6:00 PM dinner in Boston Family Members and Arrivals: - Sarah (Mom): Host, at home - James (Dad): Lands at BOS 1:00 PM from SF - Emily (Sister): Lands at BOS 2:30 PM from Chicago - Michael (Brother): Driving, arrives 3:00 PM from NY - Grandma: Needs pickup from suburban Boston Cooking Requirements: - Turkey: 4 hours cooking time - Side dishes: 2 hours preparation - Someone must stay home during cooking for fire safety Transportation Constraints: - James must rent car after landing - Emily requires airport pickup - Travel times: β Home to BOS Airport: 60 min β BOS Airport to Grandmaβs: 60 min β Home to Grandmaβs: 30 min Key Requirements: - All family members at home for 6:00 PM dinner - Turkey and sides ready by dinner time - All pickups completed with available drivers - Cooking supervision maintained
5.2. Thanksgiving Dinner Problem: P6 and P9, Testing Common Sense Reasoning and Context Management & Validation
ProblemΒ P6 considers a Thanksgiving dinner scenario in which a family of five must return to their home in a Boston suburb for a 6Β p.m.Β dinner. The problem involves coordinating departure times, managing travel logistics (including possible traffic delays), and ensuring timely arrival. TableΒ 3 formalizes these challenges as a sequential planning problem. This scenario also lays the foundation for a more advanced disruption case, which has proven difficult for standalone LLMs, as discussed inΒ P9.
5.2.1. Common Sense Augmentation
FigureΒ 4 presents a feasible schedule planned by Claude 3.7. Similarly, GPT-4o was able to generate a viable plan to ensure dinner was started on time (figure is similar and therefore not shown). However, a subtle, yet important consideration that humans typically account forβbut LLMs initially overlookedβis the time required for passengers to retrieve their luggage after landing. In practice, this process typically takes about 30 minutes before they exit the terminal.
To address this, common-sense augmentation agent was introduced into the plan. The yellow boxes in FigureΒ 4 reflect this augmentation by introducing 30 minutes for James and Emily to exit the airport.
5.2.2. Context Narrowing
Next, we use problem P9 to illustrate the attention-narrowing problem and the importance of independent validation. Problem P9 is identical to the previous instance, except that at 1 PM, James notifies the group that his plane will land at 4 PM instead of 1 PM due to an emergency detour. FigureΒ 5 shows that Claude 3.7βs reactive planning introduces constraint violations:
Fire Safety: Sarah is scheduled to leave home at 2:30 PM, leaving the oven unattended.
Travel Time: The travel time between home and BOS should be one hour, but is scheduled to be only 30 minutes.
Side Dish Preparation: The required preparation time is 2 hours, but only 90 minutes are allocated.
Dinner Time: Dinner is now scheduled for 6:30 PM, violating the 6:00 PM constraint.
Each of these violations is perplexing, given that the constraints are explicitly stated in the context. Furthermore, after multiple iterations of reactive planning within the same thread, several constraints continue to be ignored or misinterpreted (e.g., cooking safety). This highlights a key limitation in the modelβs ability to maintain global constraint awareness over sequential planning steps due to attention narrowing.
When tested with GPT-o1, all constraints were correctly observed. However, at the final step, it added 30 extra minutes to Jamesβ driving time from Boston Airport to home, citing potential traffic congestion. This kind of βclevernessβ is concerning, as an LLM may inject its own opinions at unpredictable stages in unpredictable ways. In this case, we had to explicitly prompt GPT-o1 to disclose its βcommon senseβ assumptions for verification.
This pattern suggests that during reactive planning within the same thread, the model fixates on recent adjustments while progressively disregarding earlier constraints. According to (Liu etΒ al., 2024b), some missing context may be lost from the middle of the context buffer, further contributing to systematic attention narrowing and planning inconsistencies.
Remediation: Context Management
and Reactive Planning
To remedy the problem of context narrowing/loss, employs a context management agent to checkpoint historical state transitions, unresolved dependencies, and constraints. A key design criteria is to keep the agentβs context small, to avoid its own sifferring from attention narrowing. Hence, we employ two validation agents, one for travel coordination and the other for food preparation, each keeps a context of less than .
For problem P9, we instruct the agent to store each individualβs current state, next scheduled state-transition time, and all relevant scheduling constraints.
When an unexpected event triggers reactive planning, all individuals roll back to the last saved state. The system then consolidates past and new constraints, resolving conflicts before proceeding with planning. This ensures that:
Past history is not inadvertently overridden.
New dependencies and constraints are properly integrated.
The system maintains consistency across state transitions.
By maintaining a structured history of constraint awareness, ensures robust planning, effectively mitigating LLM-driven attention narrowing and enhancing consistency in reactive temporal scheduling.
5.3. Wedding Gathering Problem: P5 and P8, Testing Transaction Property Guarantees
TableΒ 4 presents a coordination problem for travel for wedding events (Problem P5 in (Geng and Chang, 2025)). Several friends arrive at different times and locations before a 3:00Β PM wedding photo session. The challenge includes managing two vehicles for airport pick-ups (aimed at those who cannot drive or wish to cut costs) and completing critical errands, such as collecting the wedding gift and retrieving formal attire from the tailor. All activities must be scheduled to ensure that everyone arrives at the wedding venue before the photo session deadline.
Metrics: - On-time performance: Must arrive at the venue for 3:00 PM photos. Locations: Four locations: , where is Boston Airport, is Gift shop, is Tailor shop, and is Wedding venue. Travel time: (minutes) . Arrival Times: - Alex: At at 11:00 AM from Chicago (need a ride) - Jamie: At at 12:30 PM from Atlanta (need a ride) - Pat: At at 12:00 PM driving from NYC (has 5-seater car) Required Tasks: - Gift collection from (after 12:00 PM) - Clothes pickup from (by 2:00 PM) - Photos at (3:00 PM sharp) Available Resources: - One car (5-seater) with Pat, available after he is Boston - Local friend Chris (5-seater) available after 1:30 PM at Scheduling Constraints: - All tasks must complete before 3:00 PM photo time - Gift store opens at 12:00 PM - Tailor closes at 2:00 PM - Two cars must accommodate all transport needs
| Time | Activity | People |
| 11:00 AM | Alex arrives at Boston Airport (B) | Alex |
| 12:00 PM | Pat arrives at Wedding Venue (W) | Pat |
| 12:00 PM | Gift Shop (G) opens | β |
| 12:00β12:40 | Pat drives from Wedding Venue (W) to Boston Airport (B) | Pat |
| 12:30 | Jamie arrives at Boston Airport (B) | Jamie |
| 12:40β12:45 | Pat picks up Alex at Boston Airport (B) | Pat, Alex |
| 12:45β12:50 | Pat picks up Jamie at Boston Airport (B) | Pat, A., J. |
| 12:50β1:35 | Drive from BOS to Gift Shop (G) | Pat, A., J. |
| 1:30 | Chris available at Wedding Venue (W) | Chris |
| 1:35β1:50 | Collect gift at Gift Shop (G) | Pat, A., J. |
| 1:50β2:10 | Drive from Gift Shop (G) to (T) | Pat, A., J. |
| 2:00 | Tailor Shop (T) closes | β |
| 2:10β2:25 | Pick up clothes at Tailor Shop (T) | Pat, A., J. |
| 2:25β2:40 | Drive from T to Wedding Venue (W) | Pat, A., J. |
| 2:40 PM | Arrive at Wedding Venue (W) | Pat, A., J. |
| 3:00 PM | Photo session at Wedding Venue (W) | All |
5.3.1. Context Narrowing (again)
TableΒ 5 presents an infeasible schedule generated by Claude 3.7, where Pat arrives at the tailor shop (T) after closing timeβanother clear example of attention narrowing. When queried about the error, Claude 3.7 admitted that it prioritized local route optimization while losing track of global constraints.
To remedy this issue, can enforce constraint validation checkpoints at 12:50 PM, evaluating whether to send Pat to T or at 1:30 PM, when Chris becomes available to drive to T. These missed optimization opportunities can be addressed through validation protocols of βs.
In contrast, GPT-o1 correctly schedules Pat to visit the tailor shop (T) first, ensuring it is open, before proceeding to the gift shop (G) and successfully completing both errands.
However, both schedules overlook a more efficient alternative: Chris, who is available at 1:30 PM, could have handled both errands, balancing workload and improving overall efficiency. The comparative travel routes for Pat and Chris are:
Patβs route: WβB (40 min) + BβW (40 min) = 80 minutes.
Chrisβs route: WβT (15 min) + TβG (20 min) + GβW (25 min) = 60 minutes.
5.3.2. Transaction Properties
Problem P8 introduces an traffic alert:
Alert 1:00 PM: Traffic Alert, an accident near the Logan airport in Boston triples all travel times to and from the airport!
| Time | Activity | People |
| 1:00 | Traffic alert received - Pat at W | System |
| 1:05 | Pat departs W for B | Pat |
| 1:30 | Chris becomes available at W | Chris |
| 1:30 | Chris departs W for T | Chris |
| 1:45 | Chris arrives at T for clothes | Chris |
| 2:00 | Chris departs T with clothes | Chris |
| 2:15 | Chris arrives at G for gifts | Chris |
| 2:25 | Pat arrives at B (delayed by traffic) | Pat |
| 2:35 | Pat departs B with Alex & Jamie | Pat |
| 2:40 | Chris departs G with gifts | Chris |
| 2:55 | Chris arrives at W | Chris |
| 3:55 | Patβs group arrives at W (Late) | Pat |
This alert requires LLMs to replan in real time. Unfortunately, ClaudeΒ 3.7, DeepSeekΒ R1, and GPT-4o failed to react accurately to the new traffic constraints, and even GPT-o1 struggled with the precision of the planning. In contrast, can help remedy these shortcomings by maintaining both transaction state and history.
Below is a list of results from four LLMs:
@itemi
DeepSeek R1: TableΒ 6 demonstrates how DeepSeekΒ R1 fails to maintain temporal consistency in reactive planning. When the traffic alert takes effect at 1:00βPM, DeepSeek discards its execution history and attempts to create a new plan starting from that point onward. Critically, it reassigns Pat to begin driving to the airport at 1:00βPM, even though Pat had already arrived at the airport by 12:40βPM under the original schedule. This βrewriteβ of already-executed actions illustrates how LLMs can lose track of immutable past events when adapting to new conditions.
@itemi
GPT-4o: Similar to DeepSeekΒ R1, GPT-4o exhibits temporal-spatial context confusion and violates multiple constraints, demonstrating that it struggles to adapt effectively once alerts are introduced mid-plan. (Table not shown due to space limitations.)
@itemi
Claude: TableΒ 7 shows that Claude recognizes the accident, but does not update Patβs driving time from Boston Airport (departing at 12:50βPM) to the gift shop. In other words, Claude fails to fully transition into the new alert state.
@itemi
GPT-o1: TableΒ 8 shows GPT-o1βs conservative plan in which Chris handles the tailor shop, avoiding potential delays for Pat. The solution is feasible and well-founded, although it could be refined with precise spatial-temporal details (e.g., whether Pat has already passed the accident site by 1:00Β PM).
| Time | Activity | People |
| 1:00 | Traffic Alert: Accident near Airport triples travel times to/from airport | β |
| 1:00 | Current status: Pat, Alex, and Jamie en route from Airport (B) to Tailor (T) | Pat, A., J. |
| 1:00β1:10 | Emergency decision: Continue to (T) | Pat, A., J. |
| 1:10β1:25 | Arrive at (T), collect clothes | Pat, A., J. |
| 1:25β1:45 | Travel from Tailor (T) to Gift Shop (G) | Pat, A., J. |
| 1:30 | Chris is available at (W) | Chris |
| 1:30β1:45 | Chris drives from (W) to (G) | Chris |
| 1:45β2:00 | Both cars meet at (G), collect gift | P., A., J., C |
| 2:00β2:25 | Patβs car: Drive from (G) to (W) | P., A., J. |
| 2:00β2:25 | Chrisβs car: Drive from (G) to (W) | Chris |
| 2:25 | All arrive at Wedding Venue (W) | P., A., J., C. |
| 3:00 | Photo session at Wedding Venue (W) | All |
This study underscores why modern LLMs do not effectively handle disruptions: They do not keep track of fine-grained temporal-spatial state information and thus exhibit the following limitations:
State Maintenance Failure: When an alert occurs, these LLMs might discard the partial context of already-completed actions (except for GPT-o1), attempting to generate entirely new plans rather than adapt existing ones. This reveals their inability to reason about the continuous flow of time in real-world scenarios.
Temporal Consistency: They often try to modify actions that have already occurred, failing to respect the immutable nature of past events. This leads to logically infeasible schedules that effectively βrewriteβ history.
Position Tracking: They lose track of agent locations (in all cases) at critical intervals, causing plans that violate basic spatial constraints and indicating weak modeling of real-world object permanence.
Path Dependency: When conditions change mid-route, the models do not recognize that different segments of the same journey may be differently impacted by the new constraints. This demonstrates an inability to reason at a granular level about movement through time and space.
| Time | Activity | People |
| 11:00 AM | Alex arrives at Airport (B). | Alex |
| 12:00 PM | Pat departs for Airport from (W). | Pat |
| 12:30 PM | Jamie arrives at Airport (B). | Jamie |
| 12:40β12:50 PM | Pat arrives at (B), picks up Alex and Jamie; departs at 12:50 PM. | Pat, A., J. |
| 12:50β1:00 PM | Drive (B W) under normal conditions for first 10 minutes. | Pat, A., J. |
| 1:00β2:30 PM | Traffic Alert starts: remaining distance (30 min normal) becomes 90 min. Arrival at W by 2:30 PM. | Pat, Alex, Jamie |
| 1:30 PM | Chris available at (W). Departs for Tailor. | Chris |
| 1:30β1:45 | Drive (W T). | Chris |
| 1:45β1:50 | Pick up clothes at (T), closes at 2:00 PM. | Chris |
| 1:50β2:10 | Drive (T G). | Chris |
| 2:10β2:15 | Purchase gift at (G). | Chris |
| 2:15β2:40 | Drive (G W). | Chris |
| 2:30 PM | Pat, Alex, Jamie arrive at (W). | Pat, A., J. |
| 2:40 PM | Chris back at (W) with clothes and gift. | Chris |
| 3:00 PM | Wedding photo session at (W). | Everyone |
5.3.3. Remedial Approach.
By contrast, preserves both the transaction state and history, enabling it to adapt to unexpected changes without discarding vital context from past actions. Specifically, implements these strategies:
Continuous Monitoring: An active thread tracks each agentβs position and progress, allowing fine-grained analysis of how alerts impact different parts of the route.
Historical Context: All executed actions are stored as immutable facts, ensuring the system never attempts to βrewriteβ past events in the face of new constraints.
Granular Impact Analysis: recalculates individual journey segments whenever an alert occurs, accounting for how much of each route was completed pre-alert versus how much remains under the new conditions.
Feasibility Validation: Proposed modifications are checked to ensure full compliance with physical and temporal constraints, preventing invalid schedules.
Reactive Analysis:
At 1:00βPM, when the alert arrives, first checks previously executed actions. For instance, if Patβs car left the airport (B) at 12:50βPM, recognizes that 10 minutes of normal driving elapsed before the disruption. Only the remaining distance is then subjected to tripled travel times.
Impressively, GPT-o1 (TableΒ 8) is the only LLM that splits Patβs trip into pre- and post-alert segments, though it assumes Pat is definitely delayed by the accident.
formalizes this via three scenarios:
Pre-Accident Clearance: Patβs vehicle may already be past the accident site and remain unaffected.
Mid-Journey Impact: The car travels normally at first, then experiences tripled time for the remaining distance.
Direct Involvement: In the worst case, Patβs car is part of the accident, requiring major replanning or alternate transport.
By enumerating these possibilities, preserves history and integrates new alerts without overwriting executed steps, keeping schedules consistent with physical and temporal constraints.
5.4. Observations
Our experiments highlight key limitations in standalone LLMs for complex planning. The most critical is attention narrowing, where models lose sight of global constraints, miscalculate travel times, and fail to audit their plans (self-validation failure).
Common-sense gaps emerge in task handling, with models arbitrarily inserting, omitting, or misjudging time for essential steps (e.g., luggage retrieval, walking). While sometimes negligible, such errors can be critical.
LLMs also struggle with state management, often βrewritingβ past events rather than adapting, leading to inconsistent spatial-temporal reasoning. Some models lost track of agent locations, producing infeasible plans.
Our results further show varied performance across LLMs. GPT-o1 maintained partial historical awareness to some extent, though it still exhibited lapses. DeepSeekΒ R1, ClaudeΒ 3.7, and GPT-4o all showed significant stateless and context-retention issues in reactive planning scenarios.
By contrast, the framework mitigates these issues, as discussed in our conclusion.
6. Conclusion
This paper introduced , a structured multi-agent framework designed to address fundamental limitations in monolithic LLM-based planning and reasoning systems. Through systematic experimentation using the REALM benchmark, we identified and remediated four critical shortcomings in current approaches: inadequate self-validation, context narrowing, lack of transaction properties, and insufficient inter-agent coordination. Our findings demonstrate that even state-of-the-art LLMs such as Claude 3.7 and GPT-o1, despite their impressive capabilities, struggle to maintain global constraint awareness during complex planning tasks. They exhibit a tendency to fixate on recent adjustments while progressively disregarding earlier constraints, leading to inconsistent and infeasible plans. This attention narrowing effect is further exacerbated during reactive planning scenarios, where models attempt to rewrite already-executed actions rather than adapting from the current state. The framework addresses these limitations through several key innovations:
1.
Structured validation protocols: Enforce constraint checks at critical decision points.
2.
Context management mechanisms: Checkpoint historical state transitions and constraints to prevent catastrophic context loss.
3.
Transaction state preservation: Maintain immutable historical actions while enabling adaptive future planning.
4.
Specialized agent distribution: Assign focused responsibilities to prevent context overload and manage inter-agent dependency resolution.
These components work together to enable robust planning and reasoning in diverse scenarios, from complex travel logistics to time-sensitive activities with multiple interdependencies. By distributing the cognitive load among specialized agents and imposing rigorous validation, maintains consistency and adaptability even under significant disruptions.
Our findings suggest that the future of LLM-based systems lies not in scaling up monolithic models but in orchestrating smaller specialized agents through well-designed coordination mechanisms. This approach boosts performance on complex tasks while improving predictability, reliability, and transparency, all vital for mission-critical applications.
Future research will focus on mitigating context narrowing within autoregressors, developing formal verification methods for multi-agent systems, integrating with external storage for managing context (Huang etΒ al., 2020), and extending to address scientific reasoning, creative collaboration, and complex decision making under uncertainty.
Acknowledgment
I am deeply grateful to my late advisor, Hector Garcia-Molina, for his invaluable mentorship and for pioneering SAGAS, the foundational inspiration for this work.

### v2

@@
 Abstract.
-Recent LLM-based agent frameworks have demonstrated impressive capabilities in task delegation and workflow orchestration, but face significant challenges in maintaining context awareness and ensuring planning consistency.
-This paper presents , a structured multi-agent framework that addresses four fundamental limitations in current LLM approaches: inadequate self-validation, context narrowing, lacking transaction properties, and insufficient inter-agent coordination.
-By implementing specialized context management agents and validation protocols, preserves critical constraints and state information throughout complex planning processes, enabling robust and consistent decision-making even during disruptions.
-We evaluate our approach using selected problems from the REALM benchmark, focusing on sequential and reactive planning scenarios that challenge both context retention and adaptive reasoning.
-Our experiments with state-of-the-art LLMsβClaude 3.7, DeepSeek R1, GPT-4o, and GPT-o1βdemonstrate that while these models exhibit impressive reasoning capabilities, they struggle with maintaining global constraint awareness during complex planning tasks, particularly when adapting to unexpected changes.
-In contrast, the distributed cognitive architecture of shows significant improvements in planning consistency, constraint enforcement, and adaptation to disruptions in various scenarios.
+This paper introduces , a structured multi-agent architecture designed specifically to address four fundamental limitations of current LLM-based approaches: inadequate self-validation, context narrowing, absence of robust transaction-like guarantees, and insufficient inter-agent coordination.
+Unlike traditional MAS, current LLM approaches often lack critical safeguards such as transaction semantics, leading to unreliable execution and inconsistent states.
+To mitigate these challenges, integrates specialized context-management agents, compensatory rollback mechanisms, and rigorous independent validation protocols.
+Although it relaxes strict ACID constraints, particularly atomicity and isolation, adopts and adapts principles from the Saga transactional model to ensure coherent rollback and state consistency throughout complex, distributed planning processes.
+This transactional inspired approach significantly improves the robustness, constraint awareness, and adaptability of multi-agent coordination, even in the face of disruptions.
+Evaluations highlight that current standalone LLM systems, despite impressive reasoning abilities, frequently struggle with maintaining global constraints during complex planning tasks, particularly when adapting to unexpected changes.
+In contrast, the distributed transactional architecture of demonstrates significant improvements in planning consistency, constraint enforcement, and adaptive coordination during disruptions in diverse challenging scenarios.
 1. Introduction
@@
 Over the past few decades, their development has followed two primary trajectories.
-In the database community, MAS integrated robust transaction processing principles such as ACID properties (Gray, 1981), ensuring consistency and reliability for complex multi-step operations.
-In contrast, distributed systems research has prioritized coordination protocols and flexible collaboration mechanisms (Wooldridge, 2009; Durfee, 1999), enabling scalable multi-agent interactions without the constraints of strict locking or heavyweight transactional guarantees.
+In the database community, MAS traditionally integrated foundational transaction-processing principles, particularly the ACID properties (Gray, 1981; Weikum and Vossen, 2001), to ensure consistency and reliability for complex multi-step operations.
+However, for long-lived, distributed, or loosely-coupled tasks, MAS also adopted more flexible transactional models (e.g., Saga (Garcia-Molina and Salem, 1987)) to maintain robustness while relaxing strict atomicity or isolation constraints.
+In contrast, distributed systems research has prioritized coordination protocols and flexible collaboration mechanisms (Wooldridge, 2009; Durfee, 1999), enabling scalable multi-agent interactions without the constraints of strict locking or heavy-weight transactional guarantees.
 These parallel approaches have resulted in frameworks optimized for different priorities: transactional integrity versus adaptive coordination, highlighting the trade-off between strong consistency and flexible execution in practical systems.
-Recent advances in Large Language Models (LLMs) have revitalized MAS as a paradigm for sophisticated reasoning and multi-agent collaboration (Du etΒ al., 2023; Chang, 2023, 2024b).
+Recent advances in Large Language Models (LLMs) (Vaswani etΒ al., 2017; Chang etΒ al., 2024; Zhao etΒ al., 2025) have revitalized MAS as a paradigm for sophisticated reasoning and multi-agent collaboration (Du etΒ al., 2023; Chang, 2023, 2024b).
 Frameworks such as AutoGen (Wu etΒ al., 2024), LangGraph (LangChain AI, 2024), and CAMEL (Li etΒ al., 2023a) demonstrate how LLM-based agents can decompose tasks, interact across modalities, and coordinate to solve complex problems.
@@
 Compounding this issue, LLMs inherently struggle with self-validation, as highlighted by GΓΆdelβs incompleteness theorems (GΓΆdel, 1967; Chang, 2024a), making them unreliable for detecting and correcting their own errors.
-These challenges demand a transactional framework that preserves the intelligence and adaptability of LLM-based MAS while ensuring consistency and reliability in long-running, interdependent workflows.
-To address these limitations, we propose , a Transactional Multi-Agent System that extends the Saga pattern (Garcia-Molina and Salem, 1987; Weikum and Vossen, 2001)βoriginally designed for long-lived transactions by decomposing complex operations into smaller, compensable units.
-By integrating transactional logic and rollback mechanisms into LLM-based MAS, ensures that each operation is fully committed or coherently reversed, preventing partial failures and inconsistencies.
-For example, in a travel-booking scenario, a failed flight reservation would automatically trigger compensatory rollbacks of associated hotel and train bookings, maintaining consistent global states without manual intervention.
+These challenges require a transactional framework that preserves the intelligence and adaptability of LLM-based MAS while ensuring consistency and reliability in long-running, interdependent workflows.
+To address these limitations, we propose , a Transactional Multi-Agent System that extends the Saga patternβa transactional model originally developed for managing complex, long-lived transactions by decomposing them into smaller, independently committed, and compensable units.
+By integrating transactional logic and compensatory rollback mechanisms into LLM-based MAS, ensures that each individual operation within a workflow is reliably committed.
+In case of failure, clearly defined compensating transactions restore system-wide consistency.
+For example, in a travel-booking scenario, a failed flight reservation automatically triggers compensatory actions that reverse related hotel and train bookings, thus preserving consistent global states without manual intervention.
 To effectively address LLM-specific challenges and context loss problems, introduces three critical innovations that directly support our four primary contributions:
@@
 β’
-Inter-Agent Dependency Management ensures consistency by monitoring interdependencies and alerting agents to unresolved constraints, enhancing both the reliability of the system and intelligent coordination.
+Inter-Agent Dependency Management ensures consistency by monitoring inter-dependencies and alerting agents to unresolved constraints, enhancing both the reliability of the system and intelligent coordination.
 β’
-Independent Validation of Critical Junctures enhances correctness by employing specialized small-context agents with rigorous validation criteria, effectively addressing the self-validation limitations of individual LLMs.
+Independent Validation of Critical Junctures enhances correctness by employing specialized small-context agents with rigorous validation criteria, effectively addressing the limitations of self-validation of individual LLM.
 This hybrid approachβintegrating transactional processing with adaptive multi-agent intelligenceβmakes particularly effective for complex real-world applications in financial trading, healthcare, travel coordination, supply chain management, and emergency response, to name just a few.
@@
 The primary contributions of this work are as follows:
-Transactional Correctness: introduces transactional safeguards to LLM-based MAS, ensuring strong consistency across multi-agent workflows.
+Transactional Consistency via Compensation: introduces transactional safeguards and compensatory mechanisms to LLM-based MAS, ensuring reliable consistency and coherent state recovery across multi-agent workflows.
 Robust Validation: incorporates context management and verification mechanisms to improve execution correctness and prevent inconsistencies.
@@
 Transaction management has evolved significantly since the introduction of ACID properties by Gray (Gray, 1981).
-In distributed systems, the challenges of maintaining these properties across service boundaries led to alternative models such as BASE (Basically Available, Soft state, Eventually consistent) (Pritchett, 2008) and the development of specialized patterns for long-lived transactions.
-The Saga pattern, introduced by Garcia-Molina and Salem (Garcia-Molina and Salem, 1987), addresses the limitations of traditional ACID transactions in distributed, long-running processes.
-Unlike atomic transactions that rely on locks, Sagas break complex operations into sequences of smaller, local transactions, each with a corresponding compensating transaction that semantically undoes its effects.
-This approach has become fundamental in micro-service architectures and event-driven systems (Richardson, 2018), allowing them to maintain logical consistency without the performance penalties of distributed locking.
-Workflow management systems such as YAWL (vanΒ der Aalst and ter Hofstede, 2005) and more recent implementations on cloud platforms (aws, 2023; azu, 2023) have incorporated Saga-like patterns to manage complex business processes.
-However, these systems typically rely on predefined, rigid workflows with limited adaptability to unforeseen circumstancesβa limitation that becomes particularly acute when integrating with AI systems that must reason about unexpected events or novel scenarios.
-The fundamental principles of transaction management established in these systemsβatomicity, compensation-based recovery, and state consistencyβprovide a theoretical foundation for , which extends these concepts to the unique challenges of LLM-based agent workflows.
+In distributed settings, maintaining strict ACID guarantees (particularly global atomicity and isolation) across multiple independent services became impractical, leading to alternative consistency models such as BASE (basically available, soft state, eventually consistent) (Pritchett, 2008) and specialized transactional patterns tailored for long-lived transactions.
+The Saga pattern, introduced by Garcia-Molina and Salem (Garcia-Molina and Salem, 1987), explicitly addresses the limitations of traditional ACID transactions in long-lived and distributed processes.
+Unlike traditional atomic transactions relying on distributed locks, Sagas decomposes complex operations into sequences of smaller, locally atomic transactions, each paired with explicit compensating transactions that logically reverse their effects upon failure.
+This model has become fundamental to microservice architectures and event-driven systems (Richardson, 2018), enabling consistency in logical states without incurring performance penalties associated with distributed locking.
+Workflow management systems such as YAWL (vanΒ der Aalst and ter Hofstede, 2005), AWS Step Functions (aws, 2023), and Azure Logic Apps (azu, 2023) have integrated Saga-inspired transactional patterns for managing complex business processes.
+However, these implementations typically rely on predefined, rigid workflows, offering limited adaptability to unexpected events or novel scenarios, a critical limitation when integrated with AI systems that require adaptive reasoning.
+The foundational principles of local atomicity, compensatory rollback, and eventual consistency from Saga thus form a theoretical foundation for , which extends these concepts to address unique challenges in state management and consistency posed by LLM-based agent workflows.
 2.2. LLM Limitations Necessitating βs Key Requirements
-Our analysis of current LLMs reveals three fundamental limitations that directly motivate the core requirements of : the need for transactional integrity, independent validation, and strategic context preservation.
+Our analysis of current large language models (LLMs) reveals three fundamental limitations directly motivating βs core requirements: transactional integrity, independent validation, and strategic context preservation.
 Self-Validation Gap Necessitating Independent Validation
-LLMs lack mechanisms for rigorous self-validationβa limitation directly connected to GΓΆdelβs insights about formal systems (GΓΆdel, 1967).
-Recent studies confirm that while self-refinement methods (Madaan and Yazdanbakhsh, 2022; Li etΒ al., 2023b; Jiang etΒ al., 2024) iteratively improve reasoning using the LLMβs own feedback, they struggle to correct errors beyond the modelβs inherent capability ceiling (Huang etΒ al., 2024).
-In transactional contexts, this limitation manifests itself as factual inconsistencies that can lead to invalid operations, the inability to verify the feasibility of the plan, and the difficulty in maintaining consistency between operations (Yamin etΒ al., 2024).
-These challenges directly motivate βs independent validation framework.
+LLMs inherently lack robust self-validation mechanisms, a limitation stemming from intrinsic boundaries identified by GΓΆdelβs incompleteness theorems, demonstrating fundamental constraints on a systemβs ability to verify its own reasoning (GΓΆdel, 1967; Chang, 2024a).
+Recent research confirms that self-refinement techniques (Madaan and Yazdanbakhsh, 2022; Li etΒ al., 2023b; Jiang etΒ al., 2024), while iterative and beneficial, are unable to surpass inherent capability ceilings to reliably correct deeper logical errors (Huang etΒ al., 2024).
+In transactional scenarios, these validation gaps manifest as factual inconsistencies, invalid operations, and unreliable plan feasibility assessments (Yamin etΒ al., 2024).
+Thus, incorporates an independent validation framework to mitigate these inherent limitations.
 Statelessness Necessitating Transactional Integrity
-LLMs process each prompt independently, lacking inherent mechanisms to maintain the state between interactions.
-This fundamental statelessness makes transaction management essential, as there are no built-in guarantees that a sequence of operations will be treated as a cohesive unit or that failure in one operation will trigger appropriate compensations in others.
-Without transaction management, LLM-based systems risk inconsistent states, lost operations, and the inability to recover coherently from failures.
-Context Limitations Necessitating Strategic Preservation
-The self-attention mechanism of LLMs gives higher weight to immediate context, causing a narrowing of focus over longer sequences.
-Empirical measurements reveal that context retention significantly degrades beyond certain token thresholds (Xiao etΒ al., 2024; Modarressi etΒ al., 2025), and the accuracy of information recall dramatically decreases for content in middle segments of long contexts (Liu etΒ al., 2024a; Hsieh etΒ al., 2024).
-These limitations directly motivate β strategic context preservation mechanisms, as the system must explicitly track and maintain critical information that LLMs might otherwise βforgetβ during long-running transactions.
-Together, these limitations create a compelling case for addressing all three key requirements of β.
-Without comprehensive transaction management, independent validation, and strategic context preservation, LLM-based agent systems cannot provide the reliability guarantees needed for critical real-world applications.
+LLMs process each interaction independently, lacking native mechanisms to sustain state across sequential interactions.
+This fundamental statelessness necessitates explicit transactional integrity management to maintain coherent operation sequences and ensure robust failure recovery.
+Without systematic transaction management, LLM-based systems risk state inconsistency, operation losses, and incoherent recovery procedures.
+Context Limitations and Strategic Preservation
+The self attention mechanism in LLMs inherently prioritizes immediate context, significantly narrowing their effective attention windows for longer sequences.
+Empirical studies indicate substantial degradation in context retention beyond certain token thresholds (Xiao etΒ al., 2024; Modarressi etΒ al., 2025), particularly highlighting poor recall for information positioned within mid-context segments (Liu etΒ al., 2024a; Hsieh etΒ al., 2024).
+These limitations severely impact performance in extended chain-of-thought reasoning tasks (Wei etΒ al., 2022; Stechly etΒ al., 2024), where critical earlier results become susceptible to forgetting.
+Consequently, strategically preserves essential context, explicitly tracking critical information vulnerable to loss during prolonged transactions.
+Collectively, these limitations provide strong motivation for addressing all three key requirements within .
+Comprehensive transaction management, independent validation, and strategic context preservation are essential for reliably deploying LLM-based multi-agent systems (MAS) in critical real-world applications.
 2.3. Multi-Agent LLM Frameworks and Their Transaction Limitations
-Recent frameworks such as AutoGen (Wu etΒ al., 2024), LangGraph (LangChain AI, 2024), CAMEL (Li etΒ al., 2023a), and MACI (Chang, 2024a) extend multi-agent principles to LLM-based coordination but fall short in addressing the three key requirements that targets.
+Recent frameworks such as AutoGen (Wu etΒ al., 2024), LangGraph (LangChain AI, 2024), CAMEL (Li etΒ al., 2023a), and MACI (Chang, 2024a) have extended multi-agent principles to LLM-based coordination but do not fully address the three key requirements targeted by .
 Transactional Integrity Gaps
-LangGraph provides basic state management and workflow orchestration but does not implement comprehensive transaction support.
-While its graph-based workflow model allows for structured agent interactions, it lacks built-in mechanisms for atomicity guarantees, compensation handling, or failure recovery across complex multi-step operations.
-AutoGen similarly focuses on flexible agent communication but provides no systematic approach to rolling back previous actions when operations fail mid-workflow, potentially leaving the system in inconsistent states.
+LangGraph offers basic state management and workflow orchestration but lacks comprehensive transaction support.
+Although its graph-based workflow enables structured agent interactions, it does not provide built-in atomicity guarantees, systematic compensation mechanisms, or robust failure recovery.
+Similarly, AutoGen emphasizes flexible agent interactions without systematically managing compensatory actions, potentially leaving systems in inconsistent states after partial failures.
 Validation Limitations
-Current frameworks typically rely on the LLMβs inherent capabilities for validation, without the independent verification mechanisms that implements.
-This reliance on self-validation makes these systems vulnerable to the very limitations discussed in Section 2.2, as they lack robust safeguards against LLM errors, hallucinations, or inconsistencies between agent boundaries.
+Current frameworks typically rely on LLMsβ inherent self-validation capabilities, lacking independent validation procedures like those proposed in .
+This dependency exposes them to risks such as reasoning errors, hallucinations, and inconsistencies between agents, as previously discussed in SectionΒ 2.2.
 Context Preservation Challenges
-While frameworks like CAMEL maintain conversation history, they provide minimal support for strategic context preservation across long-running operations.
-Their approaches to context management focus on communication rather than the systematic tracking of operation history, dependencies, and state changes needed for reliable transaction processing.
-These frameworks serve primarily as integration platforms rather than addressing the deeper challenges related to transaction management, independent validation, and context preservation.
-Their focus on orchestration without these critical capabilities creates the gap that addresses.
+Frameworks like CAMEL maintain conversational history but offer limited support for preserving critical state transitions, inter-agent dependencies, and compensatory history.
+Their context management primarily facilitates agent communication rather than transactional reliability.
+Collectively, these frameworks serve primarily as integration platforms, failing to address comprehensive transaction management, independent validation, and strategic context preservation, critical capabilities provided directly by .
 2.4. Transaction-Oriented Approaches in LLMs
-Recent approaches have begun to address aspects of transaction management in LLM-based systems, though none provides the comprehensive solution that offers.
+Recent work has started addressing transactional management in LLM-based systems, but existing solutions remain incomplete compared to .
 Transactional Enhancements to LLM Frameworks
-Recent work on enhancing MAS with transactional capabilities (Gao etΒ al., 2024) represents an important step toward more reliable LLM workflows.
-However, these improvements focus primarily on state management without comprehensive compensation mechanisms, independent validation, or preservation of the strategic context provided by .
-Similarly, experimental transaction support in other frameworks often treats transactions as an add-on rather than a foundational component integrated throughout the system.
+Recent transactional enhancements to MAS, such as AgentScope (Gao etΒ al., 2024), primarily emphasize state management or limited rollback support without comprehensive, workflow-wide compensatory transactions or systematic independent validation as provided by .
+Experimental transactional features in other frameworks typically serve as optional add-ons rather than foundational, integrated components.
 Reliability-Focused Planning Systems
-Systems like PLASMA (Brahman etΒ al., 2024) address reliability in planning but lack transaction-specific mechanisms like compensation handling.
-Similarly, search-based approaches such as LLM-MCTS (Zhao etΒ al., 2023) and Tree-of-Thought methods (Yao etΒ al., 2024) improve plan generation but provide limited support for execution guarantees or recovery when operations fail.
-These systems typically focus on pre-execution planning rather than the runtime transaction management needed for maintaining consistency through a workflowβs lifecycle.
+Systems like PLASMA (Brahman etΒ al., 2024) focus on reliability but lack transaction-specific compensation mechanisms.
+Similarly, planning methods like LLM-MCTS (Zhao etΒ al., 2023) and Tree-of-Thought (Yao etΒ al., 2024) primarily target pre-execution reasoning, providing minimal support for transaction recovery or runtime consistency.
+Various planning methods have been proposed and (Wei etΒ al., 2025) argues that most focus only on specific problems.
 Agent Coordination with Limited Transaction Support
-Systems like ADAS and AFlow (Hu etΒ al., 2024; Zhang etΒ al., 2024) implement agent coordination protocols but lack robust transaction management.
-Their approaches to handling failures are typically ad hoc rather than following established transaction patterns like the Saga approach.
-This gap becomes particularly apparent when dealing with partial failures in complex workflows, where these systems struggle to maintain consistency across agent boundaries.
-advances beyond these approaches by providing a comprehensive framework that integrates transaction management principles, independent validation mechanisms, and strategic context preservation.
-By addressing all three key requirements simultaneously, enables LLM-based multi-agent systems to achieve reliability guarantees.
+Agent coordination frameworks like ADAS and AFlow (Hu etΒ al., 2024; Zhang etΒ al., 2024) utilize ad-hoc approaches to failure handling without robust transactional management.
+Their inability to systematically address partial failures highlights critical gaps in maintaining consistency for complex workflows.
+In contrast, integrates compensatory transaction management, independent validation, and strategic context preservation.
+By comprehensively addressing these critical requirements, significantly improves consistency, reliability, and robustness in LLM-based multi-agent workflows.
 3. System Requirements for
@@
 3.1. Transactional Requirements
-should provide transactional guarantees (Gray, 1981) that extend beyond individual agent operations to encompass entire workflows when strong consistency is required.
-For workflows requiring such guarantees, the system ensures atomicity, treating a sequence of operations performed by multiple agents as an indivisible unit.
-Given a system state , applying to results in either state (all operations applied) or state (no operations applied), with no intermediate states persisting.
-However, certain multi-agent workflows may operate effectively with eventual consistency or other relaxed guarantees when appropriate for their domain requirements.
-Consistency requirements specify that the system should transition from one valid state to another.
-Given a set of invariants , if satisfies , then should also satisfy .
-This is particularly challenging when invariants span multiple agentsβ knowledge domains.
-For isolation, concurrent workflows should not interfere with each other.
-Any two concurrent transaction sequences and should result in a state equivalent to either applying followed by , or followed by .
-Durability aims to ensure that once a workflow is committed, its effects persist despite system failures, with all state changes retrievable after system restart.
-Furthermore, while several transaction management patterns could be applied to LLM agent workflows, adopts the Saga pattern as defined by Garcia-Molina and Salem (Garcia-Molina and Salem, 1987) due to its particular suitability for long-running, distributed operations involving autonomous agents that may require compensating actions.
-maps each operation to a transaction with a corresponding compensating transaction , forming:
-where each is a local transaction and each is its corresponding compensating transaction. If any fails, all previously executed transactions are compensated by executing in reverse order.
-Alternative approaches such as two-phase distributed commit protocols (Bernstein etΒ al., 1987; Gray and Reuter, 1993) could provide stronger consistency guarantees, but would impose higher coordination overhead and potentially block agent operations during uncertainty periods.
-Similarly, eventual consistency models (Pritchett, 2008) might offer greater scalability but would sacrifice the clear semantic guarantees needed for critical workflows.
-The Saga pattern strikes an effective balance, providing meaningful transactional guarantees while accommodating the distributed and long-running nature of LLM agent operations.
+provides transactional guarantees (Gray, 1981) tailored to multi-agent workflows, extending traditional transaction semantics across agent boundaries when strong consistency is essential.
+For workflows requiring transactional coherence, the system treats sequences of operations performed by multiple agents as logically cohesive units.
+Each individual operation within is locally atomic, and in the event of a failure, compensating actions restore system-wide consistency.
+Thus, applying to a state of the system results in a fully committed new state or a coherently compensated return to the original state , preventing partial or inconsistent results.
+Consistency ensures that the system transitions between valid states, maintaining defined invariants ().
+If state satisfies , then any resulting state must also satisfy .
+This requirement is particularly challenging when invariants span multiple agentsβ knowledge domains.
+Isolation ensures that concurrent workflows execute without interference, resulting in a final state equivalent to sequentially executing workflows in some order.
+Durability guarantees that once a workflow completes, its state changes persist reliably, surviving system failures, and enabling recovery upon restart.
+Although multiple transactional patterns can support LLM agent workflows, specifically adopts the Saga pattern as defined by Garcia-Molina and Salem (Garcia-Molina and Salem, 1987), due to its suitability for long-running distributed operations requiring autonomous agent interactions and compensatory actions.
+thus maps each operation to a local transaction paired with a corresponding compensating transaction , forming:
+where a failure in any transaction triggers the execution of compensating transactions in reverse sequence.
+Alternative transaction patterns, such as two-phase commit protocols (Bernstein etΒ al., 1987; Gray and Reuter, 1993), offer stronger consistency but incur higher coordination overhead and risk agent blocking during uncertain states.
+Eventual consistency models (Pritchett, 2008) scale better but sacrifice explicit consistency guarantees essential for critical workflows.
+The Saga pattern balances meaningful transactional guarantees with the flexibility required by long-lived distributed agent operations.
 | Mechanism | Information Recorded |
@@
 3.2. Transaction State Management
-addresses the challenge of enforcing ACID guarantees in a distributed LLM agent environment with stateless agents.
-implements spatial-temporal state identification through a comprehensive transaction state management approach that preserves selective context information throughout multi-agent workflows.
-Using the MACI algorithm (Chang, 2024a) that we have previously developed, the states and dependencies of a given application specification can be automatically extracted to enable effective transaction state management.
+addresses the challenge of ensuring transactional coherence in distributed LLM agent environments characterized by stateless operations and partial context loss.
+While classical ACID properties (atomicity, consistency, isolation, durability) are too restrictive for distributed, long-running agent workflows, adopts and adapts the Saga transactional pattern.
+This pattern emphasizes local atomicity, compensating transactions, and eventual consistency, effectively managing distributed multi-agent operations without imposing strict global atomicity or isolation constraints.
+By implementing a comprehensive spatial-temporal state identification approach, selectively preserves context across multi-agent workflows, maintaining robust consistency and recoverability.
+Leveraging our previously developed MACI algorithm (Chang, 2024a), the system automatically extracts relevant states and dependencies for effective transaction management.
 3.2.1. State Dimensions and Context Preservation
-maintains three critical state dimensions to support both transaction management and context preservation:
+maintains three essential state dimensions for transaction management and context preservation:
 β’
-Application State (): The application-domain dependent data being manipulated (e.g., travel booking details).
+Application State (): Domain-specific data manipulated within transactions (e.g., travel booking details).
 β’
-Operation State (): Records of executed operations, including inputs, outputs, and effects on domain entities.
+Operation State (): Logs of executed operations, including inputs, outputs, execution status, and decision reasoning.
 β’
-Dependency State (): An explicit mapping of inter-operation dependencies (constraints) and their satisfaction criteria.
-By structuring operations, dependencies, and application data, ensures that essential information remains accessible even as LLMsβ attention mechanisms might otherwise cause older context to be lost.
-TableΒ 1 details the specific information recorded and mechanisms used to implement these state dimensions.
-Note that the decision reasoning recorded under is a critical component where LLMs make significant contributions that traditional workflow engines cannot support.
-For example, in the Thanksgiving dinner planning scenario from SectionΒ 5.2, would maintain application state for each family memberβs location and activities, operation state for travel arrangements and food preparation steps, and dependency state capturing constraints like βturkey must be in the oven by 2:00 PMβ and βJames must pick up Emily after exiting the terminal.β
+Dependency State (): Explicit mappings of inter-operation constraints and satisfaction criteria.
+By systematically structuring these state dimensions, ensures critical context remains accessible despite the inherent limitations of LLM attention mechanisms.
+TableΒ 1 details the information captured by each state dimension.
+Notably, the decision reasoning recorded under represents a key contribution from LLMs, providing detailed reasoning chains and justifications not available in traditional workflow engines.
+For example, in the Thanksgiving dinner planning scenario from SectionΒ 5.2, maintains application states tracking each personβs location, operation states for travel arrangements, and dependency states detailing constraints like βturkey must be in the oven by 2:00 PMβ and βoven must be watched for fire safety.β
 3.2.2. Dependency Tracking for Compensation
-maintains formal representations of operation dependencies in a directed graph structure:
-where represents the condition that must be satisfied for the dependency to be resolved. These conditions can be complex boolean expressions combining multiple upstream requirements, such as:
-where represents an arbitrary Boolean function that determines how the upstream conditions () are combined.
-This dependency graph is crucial for compensation planning during failures.
-When an operation fails, traverses the dependency graph to identify the operations that may be affected, creating a precise compensation sequence that ensures that the consistency of the system is maintained.
-The graph also enables to identify and preserve only the critical dependencies needed for potential compensation, reducing overhead.
+formally represents operation dependencies as a directed graph:
+where specifies the conditions necessary for dependency resolution. These conditions can be complex Boolean expressions combining multiple upstream constraints:
+with as an arbitrary Boolean operator.
+This dependency graph is essential for precise compensation planning upon transaction failures.
+When a failure occurs, traverses this graph to determine the exact sequence of compensating transactions necessary to efficiently restore global consistency.
 In the wedding coordination example from SectionΒ 5.3, when the traffic alert is received, the dependency graph allows to precisely determine which segments of Patβs journey are affected and what downstream activities (gift shop, tailor visit) need to be replanned or compensated.
 3.2.3. Intra-Agent Critical Context Tracking
-preserves critical context that LLMs might otherwise forget during extended operations.
-This includes agent specifications, intermediate reasoning states (chain-of-thought), decision justifications, and the context required for potential compensation actions.
-Empirical study (Xiao etΒ al., 2024; Modarressi etΒ al., 2025; Liu etΒ al., 2024a; Hsieh etΒ al., 2024) shows that LLMs can forget simple facts explicitly provided in context, such as travel times between locations.
-The system implements a validation cycle where agent operations produce selected logged context that undergoes validation against system specifications.
-Valid operations proceed to the next workflow node, while invalid ones trigger compensating actions before retrying.
-This ensures that:
+explicitly preserves critical contextual information typically forgotten by LLMs during extended interactions.
+The preserved context includes agent specifications, intermediate reasoning states, decision justifications, and prerequisites for compensation actions.
+Empirical studies (Xiao etΒ al., 2024; Modarressi etΒ al., 2025; Liu etΒ al., 2024a; Hsieh etΒ al., 2024) confirm that LLMs struggle to retain even explicitly provided facts over long interactions, making this context preservation vital.
+enforces a structured validation cycle in which operations produce a logged context that is independently validated.
+Validated operations proceed; failures trigger compensations using preserved historical context, ensuring that:
 β’
-Compensation mechanisms have necessary historical context.
+Compensation has sufficient historical context.
 β’
-Operations remain consistent with defined specifications.
+Operations adhere strictly to specified constraints.
 β’
-Dependency relationships maintain integrity throughout the workflow.
-This context tracking directly addresses the context narrowing issue observed in our experiments in SectionΒ 5.2, where models like Claude 3.7 lost track of critical constraints such as fire safety and travel times during reactive planning.
+Dependencies maintain integrity throughout the workflow.
+This context tracking directly addresses observed context narrowing issues from our experiments (SectionΒ 5.2), where models like Claude 3.7 failed to maintain constraints such as fire safety or travel logistics during reactive replanning.
 3.2.4. Inter-Agent Communication Protocol
-implements a structured communication protocol between agents that explicitly manages what information is passed between them. This protocol ensures that:
+defines a structured communication protocol explicitly managing information exchanged between agents. The protocol ensures:
 β’
-Each transaction agent receives precisely the information needed for its operation.
+Each transaction agent receives exactly the information necessary for the execution of the operation.
 β’
-Dependencies and constraints are properly communicated.
+Dependencies and constraints are clearly communicated.
 β’
-boundaries are respected in multi-agent operations.
-The protocol explicitly separates the agentβs working context (input prompt, reasoning space) from the transaction state.
-This separation allows agents to operate with minimal context while the system maintains complete transaction information, addressing scalability concerns in complex multi-agent workflows.
+Agent boundary constraints are consistently respected.
+This protocol separates agentsβ working contexts (prompts, reasoning chains) from transaction state management, allowing agents to function with minimal context while the broader system maintains comprehensive transaction states.
+This separation is crucial for scalability and reliability in complex multi-agent workflows.
 3.3. Independent Validation Framework
-With transaction state management ensuring persistence, we now examine how implements its independent validation of critical junctures to address the LLM self-validation limitations described in SectionΒ 1.
+With transaction state management that ensures persistence, we now detail how implements independent validation of critical junctures, directly addressing the inherent limitations of LLM self-validation discussed in SectionΒ 1.
 3.3.1. Multi-Level Validation Architecture
-implements a two-tier validation approach that separates intra-agent and inter-agent concerns.
+employs a two-tier validation strategy clearly separating intra-agent and inter-agent validation concerns.
 Intra-Agent Validation
-This validation operates within an individual LLM agent, addressing both theoretical limitations (GΓΆdelβs incompleteness) and practical challenges (attention narrowing, context forgetting).
-Syntactic Validation: Ensures outputs conform to expected formats, structures, and schemas (e.g., JSON validity).
-Semantic Validation: Verifies logical consistency, coherence, and significance of the agentβs operations.
-Factual Validation: Confirms that critical facts provided in context are maintained throughout the reasoning process (e.g., travel times, constraints).
-Constraint Adherence: Verifies that agent decisions respect all explicitly defined constraints from specifications.
-Reasoning Validation: Evaluates whether the agentβs chain-of-thought process follows logically sound inference patterns.
-Context Retention: Confirms the agent has maintained necessary contextual information required for its current operation.
-In the Thanksgiving dinner scenario, this validation would catch the fire safety violation where Sarah was scheduled to leave while the turkey was unattended in the oven.
+This validation occurs within individual LLM agents, addressing both theoretical limitations (GΓΆdelβs incompleteness) and practical challenges (attention narrowing, context loss):
+Syntactic Validation: Ensures outputs conform to expected data formats, structures, and schemas (e.g., JSON validity).
+Semantic Validation: Verifies logical coherence and meaningfulness of agent-generated outputs.
+Factual Validation: Ensures critical facts explicitly provided remain correctly preserved (e.g., travel times, constraints).
+Constraint Adherence: Confirms decisions strictly comply with all predefined system constraints.
+Reasoning Validation: Evaluates agent reasoning to ensure logically sound inference chains.
+Context Retention: Verifies that agents maintain essential contextual information throughout their operations.
+For example, intra-agent validation would detect issues like the fire-safety violation in the Thanksgiving dinner scenario, where Sarah was incorrectly scheduled to leave home while the turkey was unattended in the oven.
 Inter-Agent Validation
-This validation ensures proper information flow and dependency management across multiple agents in a workflow.
-Contract Validation: Verifies that inputs and outputs between agents conform to predefined interfaces.
-Dependency Satisfaction: Confirms that all prerequisites are met before an agent operation is executed.
-Consistency Checks: Ensures shared data elements maintain consistent meanings across agent boundaries.
-Temporal Validation: Verifies that operations happen in the correct sequence, respecting causal dependencies.
-Mutual Agreement: For collaborative decisions, confirms that multiple agents have reached consensus on shared facts.
-Transaction Boundary Integrity: Ensures that transaction ACID properties are maintained across agent operations.
-For example, in the wedding coordination scenario, inter-agent validation would ensure that Patβs position is consistently tracked across agents, preventing the error where DeepSeek R1 βrewrote historyβ by starting Patβs drive to the airport after the traffic alert despite Pat already being at the airport.
+This validation ensures correct information exchange and dependency management across multiple agents within workflows:
+Contract Validation: Confirms that inter-agent inputs and outputs align with predefined communication interfaces.
+Dependency Satisfaction: Ensures all prerequisite conditions are met before executing dependent operations.
+Consistency Checks: Validates that the shared information remains consistent and unambiguous between agent boundaries.
+Temporal Validation: Checks that the operational sequence respects causal dependencies and logical ordering.
+Mutual Agreement: For collaborative decisions, ensures consensus on critical shared facts across multiple agents.
+Transactional Integrity via Compensation: Ensures workflow-wide consistency by verifying that compensating actions correctly restore state coherence after failures.
+In the wedding coordination scenario, inter-agent validation prevents errors such as DeepSeek R1βs incorrect rewriting of Patβs position by ensuring consistent state-tracking across agent boundaries, accurately reflecting Patβs actual location after disruptions.
 3.3.2. Validation Protocols and Response Mechanisms
-To operationalize the validation types in TableΒ 2, implements structured protocols that govern system responses to validation outcomes:
-Rejection Protocol: Triggers compensating transactions upon validation failure (e.g., canceling hotel bookings if flight validation fails).
-Augmentation Protocol: Refines agent outputs while preserving agent autonomy (e.g., formatting restaurant recommendations without changing selections).
-Feedback Protocol: Captures validation results to improve future agent operations (e.g., noting that travel time estimates were consistently underestimated).
+The validation types outlined in TableΒ 2 are systematically managed by through structured protocols that define responses to validation outcomes.
+Rejection Protocol: Initiates compensating transactions upon validation failures (e.g., canceling hotel reservations if flight booking validation fails), ensuring transaction-level consistency.
+Augmentation Protocol: Enhances or clarifies agent outputs without overriding agent decisions (e.g., formatting or enriching restaurant recommendations while preserving original choices).
+Feedback Protocol: Records validation results to inform and improve subsequent agent performance (e.g., tracking consistently underestimated travel times to refine future estimates).
 These protocols directly connect validation to transaction management, ensuring that validation failures trigger appropriate compensating actions while maintaining workflow integrity across multiple agents.
 3.4. Failure Handling and Recovery
-βs failure handling leverages the transaction state and dependency tracking described in the previous sections to implement comprehensive recovery:
-Operation-Level Recovery: When individual operations fail, compensating actions execute to semantically undo their effects.
-Workflow-Level Recovery: When entire workflows fail, the system orchestrates multiple compensations in the correct order based on the captured dependency graph.
-For each compensating transaction that reverses a failed transaction , the system leverages the preserved transaction state to provide the original inputs and context that received, the outputs and side effects produced, and the specific aspects of that must be undone.
-This approach ensures reliable compensation even when the original agent is no longer available or when the operation occurred many steps earlier in a complex workflow.
+βs failure handling utilizes the transaction state management and dependency tracking described in previous sections to ensure robust recovery and maintain workflow consistency.
+Two distinct recovery levels are supported:
+Operation-Level Recovery: Upon individual operation failures, compensating transactions are executed, logically reversing the operationβs effects based on preserved inputs, outputs, and context.
+item Workflow-Level Recovery: For failures that affect entire workflows, leverages the dependency graph to orchestrate multiple compensations sequentially, ensuring global state consistency.
+For each compensating transaction that reverses a failed transaction , the system uses information preserved in the transaction state, including the original inputs, outputs, intermediate reasoning, and context, to precisely determine the aspects of to compensate.
+This explicit historical context ensures reliable compensation even when original agents become unavailable or when operations occurred many steps earlier in the workflow.
+By integrating historical context, dependency tracking, and compensatory mechanisms, maintains robust and coherent transaction recovery under complex and dynamic conditions.
 4. Design with Travel Planning
@@
 Travel Planning Specification
-This application helps users plan international trips with multiple destinations and manage complex itineraries within a specified budget.
-The system leverages LLMs to generate and refine feasible travel plans, followed by a transactional booking process implemented with to ensure consistency.
+This example application helps users plan international trips with multiple destinations and manage complex itineraries within a specified budget.
+The application leverages LLMs to generate and refine feasible travel plans, followed by a transactional booking process implemented with to ensure consistency.
 4.1. User Requirements
@@
 4.2. System Workflow
-The application workflow is divided into two main phases: itinerary planning and booking execution. FigureΒ 2 depicts the work flow with phase one in gray and phase two in green color.
+The application workflow consists of two primary phases: itinerary planning and booking execution.
+FigureΒ 2 illustrates this workflow, with phase one (planning) in gray and phase two (execution) in blue.
+Red arrows indicate compensating transactions; for example, a hotel reservationβs compensating transaction is its cancellation.
 4.2.1. Phase 1: Itinerary Planning
@@
 1.1.
-LLM generates multiple feasible itineraries based on user requirements
+LLM generates multiple feasible itineraries based on user requirements.
 1.2.
-Each itinerary includes flight options, hotel suggestions, and transportation (train) between cities
+Each itinerary specifies flight options, hotel reservations, and train transportation.
 1.3.
-System estimates costs for each component and validates against budget constraints
+Costs are estimated and validated against budget.
 1.1.
@@
 2.1.
-User reviews suggested itineraries and provides feedback
+User reviews itineraries and provides feedback.
 2.2.
-LLM adjusts plans based on feedback while maintaining budget constraints
+LLM adjusts plans according to feedback and constraints.
 2.3.
-System tracks and preserves critical context (travel dates, preferences, constraints)
+System tracks and preserves essential context (dates, preferences, constraints).
 2.4.
-Process repeats until user is satisfied
+Iterations continue until user satisfaction is reached.
 2.1.
@@
 3.1.
-User selects a final itinerary
+User selects a final itinerary.
 3.2.
-System generates a detailed breakdown of all necessary bookings
+System provides a detailed list of all required bookings.
 3.3.
-Critical dependencies between bookings are identified (e.g., hotel check-in must be after flight arrival)
+Explicitly defines booking dependencies (e.g., hotel check-in after flight arrival).
 3.1.
@@
 1.
-Booking Workflow Creation:
+Booking Workflow Creation (Saga Structure):
 1.1.
-System constructs a SAGA transaction sequence based on the finalized itinerary
+Construct a Saga transaction sequence from the finalized itinerary.
 1.2.
-Defines compensation actions for each booking step
+Clearly define each forward booking transaction () and corresponding compensating action ().
 1.3.
-Establishes validation rules for each transaction
+Establish validation rules for each transaction to trigger compensations if necessary.
 1.1.
 2.
-Execution:
+Execution (Saga Forward Transactions):
 2.1.
-International Flight Booking: SFO Berlin
+: International Flight Booking (SFO Berlin)
 2.2.
-Berlin Hotel Booking
+: Berlin Hotel Booking
 2.3.
-Train Booking: Berlin Cologne
+: Train Booking (Berlin Cologne)
 2.4.
-Cologne Hotel Booking
+: Cologne Hotel Booking
 2.5.
-International Flight Booking: Cologne SFO
+: International Return Flight Booking (Cologne SFO)
 2.1.
 3.
-Exception Handling:
+Exception Handling and Compensation:
 3.1.
-Potential failures:
-Flight unavailability or price changes
-Hotel rooms no longer available
-Train scheduling conflicts
+Potential failure scenarios:
+Flights unavailable or over the budget.
+Hotel rooms no longer available or price increased.
+Train scheduling conflicts.
 3.2.
-Compensation strategies:
-Cancel previous bookings (with cancellation policies)
-Replan affected portions of the itinerary
-Suggest alternatives within similar price range
+Compensation and Replanning strategies:
+Immediate execution of compensating actions (), such as canceling previously confirmed bookings according to predefined policies.
+After compensations, replan only the affected portion of the itinerary.
+Provide alternative options within user budget constraints.
 3.1.
@@
 For our experiments, we focused on two medium-tier sequential planning challenges (problems #5 and #6) and two reactive planning challenges (problems #8 and #9).
-We evaluated four LLMsβClaudeΒ 3.7, DeepSeekΒ R1, GPT-4o, and GPT-o1βalongside our proposed framework. All experiments were conducted between MarchΒ 12 andΒ 17, 2025.
+We evaluated four LLMsβClaudeΒ 3.7 (Anthropic, 2024), DeepSeekΒ R1 (DeepSeek-AI etΒ al., 2025), GPT-4o (OpenAI, 2024), and GPT-o1βalongside our proposed framework.
+All experiments were conducted between MarchΒ 12 andΒ 17, 2025.
+The source code of for conducting these experiments is located at .
+We evaluated four LLMsβClaudeΒ 3.7 (Anthropic, 2024), DeepSeekΒ R1 (DeepSeek-AI etΒ al., 2025), GPT-4o (OpenAI, 2024), and GPT-o1βalongside our proposed framework.
+All experiments were conducted between MarchΒ 12 andΒ 17, 2025.
+The source code of for conducting these experiments is available at (Geng, 2025).
 Objective: Coordinate family arrivals and dinner preparation for 6:00 PM dinner in Boston Family Members and Arrivals: - Sarah (Mom): Host, at home - James (Dad): Lands at BOS 1:00 PM from SF - Emily (Sister): Lands at BOS 2:30 PM from Chicago - Michael (Brother): Driving, arrives 3:00 PM from NY - Grandma: Needs pickup from suburban Boston Cooking Requirements: - Turkey: 4 hours cooking time - Side dishes: 2 hours preparation - Someone must stay home during cooking for fire safety Transportation Constraints: - James must rent car after landing - Emily requires airport pickup - Travel times: β Home to BOS Airport: 60 min β BOS Airport to Grandmaβs: 60 min β Home to Grandmaβs: 30 min Key Requirements: - All family members at home for 6:00 PM dinner - Turkey and sides ready by dinner time - All pickups completed with available drivers - Cooking supervision maintained
@@
 However, at the final step, it added 30 extra minutes to Jamesβ driving time from Boston Airport to home, citing potential traffic congestion.
-This kind of βclevernessβ is concerning, as an LLM may inject its own opinions at unpredictable stages in unpredictable ways.
-In this case, we had to explicitly prompt GPT-o1 to disclose its βcommon senseβ assumptions for verification.
+This kind of βclevernessβ is on one hand appreciated because the LLM injects common sense.
+However, it is also concerning, as an LLM may inject its own opinions at unpredictable stages in unpredictable ways.
+For common sense injection, human supervision would be preferable to ensure the applied common sense actually reflects shared understanding that is truly common among all folks, particularly those living in the Boston area.
 This pattern suggests that during reactive planning within the same thread, the model fixates on recent adjustments while progressively disregarding earlier constraints.
 According to (Liu etΒ al., 2024b), some missing context may be lost from the middle of the context buffer, further contributing to systematic attention narrowing and planning inconsistencies.
-Remediation: Context Management
-and Reactive Planning
-To remedy the problem of context narrowing/loss, employs a context management agent to checkpoint historical state transitions, unresolved dependencies, and constraints.
-A key design criteria is to keep the agentβs context small, to avoid its own sifferring from attention narrowing.
-Hence, we employ two validation agents, one for travel coordination and the other for food preparation, each keeps a context of less than .
-For problem P9, we instruct the agent to store each individualβs current state, next scheduled state-transition time, and all relevant scheduling constraints.
+5.2.3. Remediation: Context Management Β and Reactive Planning
+To address the issue of context narrowing and loss, employs a context management agent to checkpoint historical state transitions, unresolved dependencies, and constraints.
+A key design criterion is to keep the agentβs context small to prevent it from suffering from attention narrowing itself.
+Hence, we employ two validation agents: one for travel coordination and another for food preparation, each maintaining a context of less than .
+The travel coordination agent records in external storage the temporal-spatial states of each individual and relevant temporal constraints.
+For problem P9, it stores the individualβs current state, next scheduled state-transition time, and all relevant constraints.
 When an unexpected event triggers reactive planning, all individuals roll back to the last saved state.
-The system then consolidates past and new constraints, resolving conflicts before proceeding with planning.
+The system then consolidates past and new constraints, resolving conflicts through βcompensationalβ schedule cancellation before proceeding with rescheduling.
 This ensures that:
-Past history is not inadvertently overridden.
-New dependencies and constraints are properly integrated.
-The system maintains consistency across state transitions.
+Past history is preserved and not inadvertently overridden.
+New dependencies and constraints are properly restored (e.g., oven safety watch) and integrated.
+Consistency across state transitions is maintained.
 By maintaining a structured history of constraint awareness, ensures robust planning, effectively mitigating LLM-driven attention narrowing and enhancing consistency in reactive temporal scheduling.
@@
 Problem P8 introduces an traffic alert:
-Alert 1:00 PM: Traffic Alert, an accident near the Logan airport in Boston triples all travel times to and from the airport!
+Alert 1:00 PM: Traffic Alert, an accident near Logan Airport in Boston triples all travel times to and from the airport! Only correctly handles this alert.
+| Time | Activity | People |
+| 1:00 | Traffic Alert: Accident near Airport triples travel times to/from airport | β |
+| 1:00 | Current status: Pat, Alex, and Jamie en route from Airport (B) to Tailor (T) | Pat, A., J. |
+| 1:00β1:10 | Emergency decision: Continue to (T) | Pat, A., J. |
+| 1:10β1:25 | Arrive at (T), collect clothes | Pat, A., J. |
+| 1:25β1:45 | Travel from Tailor (T) to Gift Shop (G) | Pat, A., J. |
+| 1:30 | Chris is available at (W) | Chris |
+| 1:30β1:45 | Chris drives from (W) to (G) | Chris |
+| 1:45β2:00 | Both cars meet at (G), collect gift | P., A., J., C |
+| 2:00β2:25 | Patβs car: Drive from (G) to (W) | P., A., J. |
+| 2:00β2:25 | Chrisβ car: Drive from (G) to (W) | Chris |
+| 2:25 | All arrive at Wedding Venue (W) | P., A., J., C. |
+| 3:00 | Photo session at Wedding Venue (W) | All |
 | Time | Activity | People |
@@
 In contrast, can help remedy these shortcomings by maintaining both transaction state and history.
-Below is a list of results from four LLMs:
+The following is a list of results from four LLMs:
 @itemi
-DeepSeek R1: TableΒ 6 demonstrates how DeepSeekΒ R1 fails to maintain temporal consistency in reactive planning.
+Claude: TableΒ 6 shows that Claude 3.7 recognizes the accident, but does not update Patβs driving time from Boston Airport (departing at 12:50βPM) to the gift shop.
+In other words, Claude 3.7 fails to fully transition into the new alert state.
+@itemi
+DeepSeek R1: TableΒ 7 demonstrates how DeepSeekΒ R1 fails to maintain temporal consistency in reactive planning.
 When the traffic alert takes effect at 1:00βPM, DeepSeek discards its execution history and attempts to create a new plan starting from that point onward.
@@
 @itemi
-Claude: TableΒ 7 shows that Claude recognizes the accident, but does not update Patβs driving time from Boston Airport (departing at 12:50βPM) to the gift shop.
-In other words, Claude fails to fully transition into the new alert state.
-@itemi
 GPT-o1: TableΒ 8 shows GPT-o1βs conservative plan in which Chris handles the tailor shop, avoiding potential delays for Pat.
-The solution is feasible and well-founded, although it could be refined with precise spatial-temporal details (e.g., whether Pat has already passed the accident site by 1:00Β PM).
-| Time | Activity | People |
-| 1:00 | Traffic Alert: Accident near Airport triples travel times to/from airport | β |
-| 1:00 | Current status: Pat, Alex, and Jamie en route from Airport (B) to Tailor (T) | Pat, A., J. |
-| 1:00β1:10 | Emergency decision: Continue to (T) | Pat, A., J. |
-| 1:10β1:25 | Arrive at (T), collect clothes | Pat, A., J. |
-| 1:25β1:45 | Travel from Tailor (T) to Gift Shop (G) | Pat, A., J. |
-| 1:30 | Chris is available at (W) | Chris |
-| 1:30β1:45 | Chris drives from (W) to (G) | Chris |
-| 1:45β2:00 | Both cars meet at (G), collect gift | P., A., J., C |
-| 2:00β2:25 | Patβs car: Drive from (G) to (W) | P., A., J. |
-| 2:00β2:25 | Chrisβs car: Drive from (G) to (W) | Chris |
-| 2:25 | All arrive at Wedding Venue (W) | P., A., J., C. |
-| 3:00 | Photo session at Wedding Venue (W) | All |
-This study underscores why modern LLMs do not effectively handle disruptions: They do not keep track of fine-grained temporal-spatial state information and thus exhibit the following limitations:
-State Maintenance Failure: When an alert occurs, these LLMs might discard the partial context of already-completed actions (except for GPT-o1), attempting to generate entirely new plans rather than adapt existing ones.
-This reveals their inability to reason about the continuous flow of time in real-world scenarios.
-Temporal Consistency: They often try to modify actions that have already occurred, failing to respect the immutable nature of past events.
-This leads to logically infeasible schedules that effectively βrewriteβ history.
-Position Tracking: They lose track of agent locations (in all cases) at critical intervals, causing plans that violate basic spatial constraints and indicating weak modeling of real-world object permanence.
-Path Dependency: When conditions change mid-route, the models do not recognize that different segments of the same journey may be differently impacted by the new constraints.
-This demonstrates an inability to reason at a granular level about movement through time and space.
+The solution is feasible but coarse-grained, as it doesnβt leverage precise spatial-temporal reasoning about Patβs current position relative to the accident location.
+A more refined approach would first determine whether Pat has already passed the accident site by 1:00Β PM, which could eliminate unnecessary detours and resource reallocations.
+This highlights the difference between merely finding a feasible solution versus optimizing based on detailed state information.
 | Time | Activity | People |
@@
 | 3:00 PM | Wedding photo session at (W). | Everyone |
-5.3.3. Remedial Approach.
-By contrast, preserves both the transaction state and history, enabling it to adapt to unexpected changes without discarding vital context from past actions. Specifically, implements these strategies:
-Continuous Monitoring: An active thread tracks each agentβs position and progress, allowing fine-grained analysis of how alerts impact different parts of the route.
-Historical Context: All executed actions are stored as immutable facts, ensuring the system never attempts to βrewriteβ past events in the face of new constraints.
-Granular Impact Analysis: recalculates individual journey segments whenever an alert occurs, accounting for how much of each route was completed pre-alert versus how much remains under the new conditions.
-Feasibility Validation: Proposed modifications are checked to ensure full compliance with physical and temporal constraints, preventing invalid schedules.
-Reactive Analysis:
-At 1:00βPM, when the alert arrives, first checks previously executed actions.
-For instance, if Patβs car left the airport (B) at 12:50βPM, recognizes that 10 minutes of normal driving elapsed before the disruption.
-Only the remaining distance is then subjected to tripled travel times.
-Impressively, GPT-o1 (TableΒ 8) is the only LLM that splits Patβs trip into pre- and post-alert segments, though it assumes Pat is definitely delayed by the accident.
-formalizes this via three scenarios:
-Pre-Accident Clearance: Patβs vehicle may already be past the accident site and remain unaffected.
-Mid-Journey Impact: The car travels normally at first, then experiences tripled time for the remaining distance.
-Direct Involvement: In the worst case, Patβs car is part of the accident, requiring major replanning or alternate transport.
-By enumerating these possibilities, preserves history and integrates new alerts without overwriting executed steps, keeping schedules consistent with physical and temporal constraints.
+5.3.3. LLM Limitations and Remediation
+This study reveals critical limitations in how modern LLMs handle disruptions in planning scenarios:
+State Maintenance Failure: When an alert occurs, these LLMs might discard the partial context of already-completed actions, attempting to generate entirely new plans rather than adapt existing ones.
+This reveals their inability to reason about the continuous flow of time in real-world scenarios.
+Temporal Inconsistency: They attempt to modify immutable past events.
+Position Tracking: Agent locations are lost at critical intervals.
+Path Dependency: Models cannot recognize that different segments of a journey may be differently impacted by an alert.
+By contrast, implements a comprehensive remediation approach through fine-grained compensation:
+Persistent Context Repository: maintains an external state repository that captures the complete world state at each checkpoint, enabling reliable rollback and forward projection regardless of attention constraints in the planning agent.
+Immutable Action Logging: All executed actions are recorded as immutable transactions in a persistent log, ensuring that historical events remain consistent even when replanning occurs, preventing the βamnesia effectβ common in LLM planners.
+Compensatory Planning: When disruptions occur, doesnβt simply replan from scratch but applies compensatory actions specifically designed to address the deviation while preserving as much of the original plan as possible.
+Constraint Consistency Validation: The system continuously validates that new plans remain consistent with both physical limitations and temporal dependencies established in earlier planning phases.
+Compensatory Analysis:
+When faced with disruptions (e.g., the 1:00 PM traffic alert in our wedding scenario), executes a structured compensation process:
+This approach enables three key capabilities: (1) partial journey compensation for route segments, (2) strategic resource reallocation when needed, and (3) principled constraint relaxation with appropriate compensatory actions.
+Here, the key state to facilitate a precise resolution is to answer the question: βHas Patβs vehicle passed the accident location at 1:00 PM (and hence unaffected)?β The answer determines the remaining time required to reach the originally scheduled destination, the Tailor.
+In such case, no rescheduling is required.
+If Patβs car is unfortunately involved in the accident, a more comprehensive replanning approach would be necessary to accommodate this significant disruption.
 5.4. Observations
-Our experiments highlight key limitations in standalone LLMs for complex planning.
-The most critical is attention narrowing, where models lose sight of global constraints, miscalculate travel times, and fail to audit their plans (self-validation failure).
-Common-sense gaps emerge in task handling, with models arbitrarily inserting, omitting, or misjudging time for essential steps (e.g., luggage retrieval, walking).
-While sometimes negligible, such errors can be critical.
-LLMs also struggle with state management, often βrewritingβ past events rather than adapting, leading to inconsistent spatial-temporal reasoning.
-Some models lost track of agent locations, producing infeasible plans.
-Our results further show varied performance across LLMs.
-GPT-o1 maintained partial historical awareness to some extent, though it still exhibited lapses.
-DeepSeekΒ R1, ClaudeΒ 3.7, and GPT-4o all showed significant stateless and context-retention issues in reactive planning scenarios.
-By contrast, the framework mitigates these issues, as discussed in our conclusion.
+Our experiments across multiple LLMs (GPT-o1, DeepSeekΒ R1, ClaudeΒ 3.7, GPT-4o) highlight consistent limitations in complex planning scenarios.
+While GPT-o1 showed partial historical awareness, all models exhibited attention narrowing, self-validation failure, and inconsistent spatial-temporal reasoning.
+TableΒ 9 summarizes βs context management and compensation mechanisms directly address these limitations of LLMs.
+| Capability | Standard LLMs | |
+| Maintains historical actions | Partial/None | Full |
+| Partial journey compensation | Rarely | Always |
+| Constraint consistency checking | Ad-hoc | Systematic |
+| Handles attention narrowing | Vulnerable | Resistant |
+| Physical-temporal consistency | Inconsistent | Guaranteed |
 6. Conclusion
-This paper introduced , a structured multi-agent framework designed to address fundamental limitations in monolithic LLM-based planning and reasoning systems.
-Through systematic experimentation using the REALM benchmark, we identified and remediated four critical shortcomings in current approaches: inadequate self-validation, context narrowing, lack of transaction properties, and insufficient inter-agent coordination.
-Our findings demonstrate that even state-of-the-art LLMs such as Claude 3.7 and GPT-o1, despite their impressive capabilities, struggle to maintain global constraint awareness during complex planning tasks.
-They exhibit a tendency to fixate on recent adjustments while progressively disregarding earlier constraints, leading to inconsistent and infeasible plans.
-This attention narrowing effect is further exacerbated during reactive planning scenarios, where models attempt to rewrite already-executed actions rather than adapting from the current state.
-The framework addresses these limitations through several key innovations:
-1.
-Structured validation protocols: Enforce constraint checks at critical decision points.
-2.
-Context management mechanisms: Checkpoint historical state transitions and constraints to prevent catastrophic context loss.
-3.
-Transaction state preservation: Maintain immutable historical actions while enabling adaptive future planning.
-4.
-Specialized agent distribution: Assign focused responsibilities to prevent context overload and manage inter-agent dependency resolution.
-These components work together to enable robust planning and reasoning in diverse scenarios, from complex travel logistics to time-sensitive activities with multiple interdependencies.
-By distributing the cognitive load among specialized agents and imposing rigorous validation, maintains consistency and adaptability even under significant disruptions.
-Our findings suggest that the future of LLM-based systems lies not in scaling up monolithic models but in orchestrating smaller specialized agents through well-designed coordination mechanisms.
-This approach boosts performance on complex tasks while improving predictability, reliability, and transparency, all vital for mission-critical applications.
-Future research will focus on mitigating context narrowing within autoregressors, developing formal verification methods for multi-agent systems, integrating with external storage for managing context (Huang etΒ al., 2020), and extending to address scientific reasoning, creative collaboration, and complex decision making under uncertainty.
+We introduced , a structured transactional multi-agent framework specifically designed to overcome fundamental weaknesses of monolithic LLM-based planning systems.
+Through systematic experimentation using the REALM benchmark, we identified and addressed four critical shortcomings of existing approaches: inadequate self-validation, context narrowing, absence of transaction properties, and insufficient inter-agent coordination.
+Our results demonstrate that even state-of-the-art LLMs, such as Claude 3.7 and GPT-o1, despite strong reasoning capabilities, often fixate on recent context while neglecting critical earlier constraints, leading to inconsistent and infeasible plans.
+This issue is particularly pronounced in reactive planning scenarios, where models frequently attempt to retroactively rewrite past actions rather than adapting from the current state.
+The framework addresses these limitations through four key innovations.
+Structured validation protocols: Independent constraint checks to overcome inadequate self-validation.
+Context management mechanisms: Strategic checkpointing to mitigate context narrowing in long sequences.
+Transactional state preservation: Immutable historical records and compensatory mechanisms to address missing transaction properties.
+Specialized agent distribution: Explicit role specialization and dependency tracking to resolve insufficient inter-agent coordination.
+Collectively, these innovations enable robust and coherent planning in diverse real-world scenarios, from complex travel logistics to dynamic, time-sensitive tasks.
+By delegating tasks among specialized agents and enforcing rigorous transactional validation, significantly improves consistency, reliability, transparency, and adaptabilityβqualities essential for mission-critical applications.
+Future research will address intrinsic context narrowing in autoregressive models, develop formal verification methods for transactional multi-agent systems, integrate advanced external context-management tools (Huang etΒ al., 2020), and extend to complex domains such as scientific reasoning, creative collaboration, and decision-making under uncertainty.
 Acknowledgment
 I am deeply grateful to my late advisor, Hector Garcia-Molina, for his invaluable mentorship and for pioneering SAGAS, the foundational inspiration for this work.
+Supplementary Materials
+Benchmark: Our established benchmark (Geng and Chang, 2025) for evaluating multi-agent systems is located at
+https://github.com/genglongling/REALM-Bench. -
+Experiment Source Code: The source code of the for solving the four benchmark problems is located at
+https://github.com/genglongling/SagaLLM.